### v3

# Digital currency hardware wallets and the essence of money

Abstract
Many proposals for the design and implementation of digital wallets assume that the purpose of the wallet is to enable offline payments via custodial accounts, ignoring the real problems faced by individuals and businesses that engage in retail payments, such as the anticompetitive behaviour of payment platforms and the decline of cash. More importantly, the proposals ignore the raison d’être of digital currency as a kind of digital money that can be held independently of custodians. Finally, the proposals demonstrate a profound lack of imagination about the nature of digital money and the devices that could be used to hold, manage, and exchange it. From these presumptions flows a set of architectural requirements that stifle the promise of digital currency to deliver novel and efficient ways to exchange value in the digital economy. In this article, we critically assess the essential problems that digital currency solutions are being proposed to solve, particularly with respect to the future of payments and the future of cash. We assess the validity of common justifications for account-based payments and certified hardware in the context of alternative designs, limitations, and trade-offs. We conclude that the interests of consumers would be better served by design approaches to digital currency that anticipate that digital assets would be held outside accounts, stored offline, but transacted online, without requiring the use of trusted hardware.
1 Introduction
The proliferation of cashless payments has given rise to a growing preponderance of merchants who refuse to accept cash, as well as the suggestion that private-sector payment systems could subvert the authority of central banks to implement monetary policy. This development has, in turn, focussed the attention of central banks and other financial institutions around the world on digital currency, particularly the idea that central banks could issue central bank digital currency (CBDC) for retail use. In principle, retail CBDC would be a direct obligation of the central bank, like cash or central bank reserves, but existing in digital form and held by ordinary consumers and non-financial businesses.
Many proponents of retail CBDC have argued that it must be “cash-like” to deliver value to end-users. This argument is at least partially right. For example, it might be expected that a cash-like system would afford users privacy by design, resistance to discrimination and profiling, and the ability to possess and control the assets that they use to engage with the economy. These are certainly desirable properties. However, some authorities have interpreted the argument that a digital payment system must be cash-like to mean that a digital payment system must enable dual-offline transactions, namely those without connectivity of any kind between either of the counterparties and any third parties, in contrast to “single-offline” transactions, wherein a mutually trusted third party is reachable by one of the counterparties. We suggest that this argument constitutes a leap of logic that must be examined carefully, starting with its implicit assumptions that any new payment system must be appropriate for all scenarios, that existing solutions such as cash are insufficient to handle payment scenarios in which neither party has access to a network connection, and that there is a specific need for future digital currency infrastructure to enable a kind of digital transaction that is not commonly available to users today.
Importantly, the idea that digital currency transactions must mirror the current system for electronic retail payments demonstrates an assumption that the main promise of digital currency concerns accessing custodial accounts whilst offline. But this assumption ignores decades of history of digital currency, wherein the most groundbreaking systems, including not only Bitcoin but also DigiCash two decades earlier, drew attention by introducing a way to hold money outside of custodial accounts. The idea that central banks would never support such an architecture ignores the fact that central banks have always supported the issuance and circulation of tokens outside accounts, in the form of cash. Within this framework, offering cash in digital form is the next logical step.
Dual-offline transactions impose additional security requirements to prevent double-spending attacks, in addition to the new security requirements applicable to digital bearer instruments in general. Some prominent experts have proposed “trusted computing” [1, 2] as a solution to the security problem, reasoning that if the security of bearer instruments is a cause for concern, then perhaps there is benefit to be realised in making an effort to ensure that the devices that can hold them must meet a certain standard. After all, many users are accustomed to carrying certified hardware, and the prevalence of certified hardware among common user devices might be seen as a justification for their use in the context of digital currency. For example, smart cards commonly rely upon tamper-resistant chips containing secrets issued by their manufacturers [3]. Mobile phones often contain secure enclaves to restrict what their users can do with their devices, for example to prevent users from copying multimedia files.
We also consider the conflation of technology, which can be freely and directly deployed by an individual or business to perform a function, with services, which are functions performed by third parties. In a world of smartphone applications for which most of the computation is nearly always performed by remote service providers, it is easy to ignore this distinction. However, since the ability to possess and control money directly is essential to its function, we must be careful to distinguish between technology that is under the control of the end-user and technology that is under the control of a third party.
Of course, the fact that certain methods are commonly used in some contexts does not justify their use in others, and it certainly does not justify mandating the use of such methods in the context of future infrastructure that intrinsically must be available for general use by the public. It is all too easy for those in the payments industry to assume that everyone has or should have a smartphone and a bank account. Mandating the use of certified hardware carries many costs, and the benefits of certified hardware are not as clear as its proponents claim.
The aforementioned considerations raise the question of how end-users will hold digital currency assets, and in particular, what the role of a wallet should be. ISO 22739:2024 defines a wallet as an “application or mechanism used to generate, manage, store or use private keys and public keys or other digital assets” [42]. The simplicity of this definition would seem to understate the potentially complex features required to implement token management and custody, an exploration of which is given by Barresi and Zatti [5]. We address some of the assumptions that Barresi and Zatti make about the design of wallets and suggest different ways of viewing the problem. In particular, we critically address the arguments that trusted computing is an indispensable enabler of digital currency solutions, that digital currency must be held in accounts, that wallets must be linked to the identity of a user, and that a wallet is properly seen as “not a place where you can store cash [but] that helps you keep track of it” [5].
The rest of the article is organised as follows. In the next section, we consider the arguments in favour of offline payments, toward an understanding of precisely what is meant by the use of the term “offline”, as well as the specific kinds of offline payments for which trusted computing has been proposed as a necessary solution and the underlying motivations for this reasoning. In Section 3, we critically examine the requirement for trusted computing in the context of central bank digital currency, identifying several key problems that undermine its perceived value. In Section 4, we argue that token-based architectures are appropriate for privacy and self-custody. In Section 5, we discuss the role of the issuer as distinct from the trusted third parties that facilitate transactions. In Section 6, we offer an alternative approach that does not rely upon trusted computing and argue that this and similar approaches are better-suited for retail digital currency infrastructure in the general case. In Section 7, we summarise our arguments.
2 The sophistry of dual-offline payments
Recently, several prominent organisations, including the International Monetary Fund [6] as well as central banks such as the Bank of England [7] and the Bank of Canada [8], have released reports suggesting that dual-offline payments are desirable or even essential to the functionality of a future retail CBDC system. Such bold statements are striking, considering the prevailing lack of consensus about what the requirements for digital currency should be in general. In this section, we shall critically explore the argument that the purpose of digital currency is to facilitate access to custodial accounts in a dual-offline payment scenario.
2.1 Kinds of payments
First, we must ask: What exactly constitutes an “offline” payment? We consider the most common kinds of retail payment mechanisms and assess them in terms of whether they are online or offline:
  • E-commerce transactions. Transactions that involve consumers using personal computers or mobile devices to interact with websites require the internet by definition. They represent a significant and growing share of the total value consumers spend in the economy, and naturally, they require consumers to be online.
  • Bank transfers. Bank transfers are mechanisms for transferring money from a sender’s account with a sending bank to a recipient’s account with a receiving bank. As account-holders, both the sender and the recipient must be identified. Bank transfers can be arranged online, or an account-holder can arrange a bank transfer via a conversation with a bank teller or a computer controlled by the bank. The sending bank debits the sender’s account, and the receiving bank credits the recipient’s account. Settlement between the banks is managed via an interbank clearing and settlement network. In most cases, interbank settlement networks cannot guarantee “immediate” settlement within a few seconds, and bank transfers are usually not considered appropriate at the point of sale or for transactions that require acceptance in real-time.
  • Electronic card payments at the point of sale (POS). Electronic card payments provide a mechanism by which merchants can accept in-person payments from consumers in real-time. The consumer’s bank (the issuer) and the merchant’s bank (the acquirer) both participate in a payment network. The issuer provides a card to the consumer, and the acquirer offers to connect a merchant’s POS device to the payment network. The card serves as a form of identification: Its purpose is to specify a particular bank account and to identify its user as the beneficiary of that account. To protect against unauthorised copying of the identification credential, cards are usually designed with features that provide tamper-resistance. The client is not assumed to have an Internet connection, and the POS device provides a way for the card to authenticate itself to the payment network. If the authentication is successful, then the payment network effectively underwrites the transaction, implicitly providing credit to the issuer and prompting the acquirer to accept the transaction immediately. The payment network informs the merchant that the transaction is approved, and the merchant can then accept the payment in real-time. Merchants pay interchange fees to support this mechanism.
  • Closed-loop payment systems at the point of sale. With closed-loop payment networks, both consumers and merchants are primary customers of a platform operator, which maintains accounts with balances for all of its customers. When a consumer pays a merchant in-person, the platform operator debits the consumer’s account and credits the merchant’s account. For a transaction to succeed, either the consumer or the merchant must have an active connection to the platform operator. The consumer must authenticate successfully to the platform operator, although this can be done via the merchant’s device, as it is with card payment networks. Crucially, banks are not generally involved in transactions; the platform operator manages the process.
  • Cash. Cash comprises a set of physical tokens, each of which is a direct obligation of a central bank or monetary authority. Cash is held directly by individuals and non-financial businesses as a store of value, and it is also held by banks to service withdrawals, wherein banks debit accounts of customers in exchange for providing them with cash. Cash is fungible, which is to say that individual cash tokens are mutually substitutable and undifferentiated in practice. Since the consummation of a cash transaction is the exchange of a physical token, cash transactions between parties are offline by definition. Regulators generally do not have a direct view into cash transactions, and assessment of regulatory and tax compliance depends upon accounting and reporting.
E-commerce transactions are certainly online, and cash transactions are certainly offline. However, whether point of sale transactions and bank transfers are online or offline is a matter of perspective. For example:
  • Is it enough for one party but not both to be connected to the Internet?
  • Is it necessary to be able to reach a particular service run by a specific operator, or is it sufficient to be able to reach any one of a set of operators?
  • Can a transaction to be provisionally allocated using the network but consummated when both parties are disconnected from the network?
Notwithstanding some measure of understandable confusion and inconsistency in the use of the term “offline” accompanying calls for considering offline payments to be an essential feature of digital currency, it can be said that some measure of online technology supports virtually all retail payments today, excepting only payments in cash, which has seen a dramatic decline in use throughout the world over the past decade.
With this in mind, it is reasonable to ask: What problem do proponents of retail CBDC intend to solve, for which dual-offline payments constitute a better solution than online payments involving assets held offline? We acknowledge that there are some limited real-world use cases for dual-offline payments, such as payments aboard transportation systems that lack reliable network connectivity throughout the system [9]. However, such examples are rare indeed, owing to the prevalence of network connectivity in locations suitable for digital transactions. Furthermore, we note that payment scenarios similar to this example do not actually require dual-offline payments; their requirements could also be satisfied by time-shifting, in which consumers purchase e-money tokens offline and consummate the transactions offline by selectively revealing proofs of purchase [10]. And so, it is fair to ask: Might the idea of dual-offline payments constitute a solution in search of a problem, an ephithet that a report by the UK House of Lords prominently ascribed to CBDC in general [11]?
2.2 Digital currency use cases
One might argue that conducting transactions using digital currency held by custodians is, from the perspective of users and counterparties, not much different than conducting transactions practice of using ordinary bank accounts, with a possible exception of a different regulatory environment for the digital currency custodian. The custodian is still a gatekeeper that could fail to meet its obligations; the user is still bound to the custodian by identification requirements and still does not hold the assets directly. One might also argue that digital currency held by custodians is even less useful than bank deposits, if it is assumed that digital currency tokens cannot be rehypothecated and therefore cannot contribute to money creation.
We might reasonably conclude that non-custodial wallets, which allow users to hold digital assets directly, are a foundational benefit of digital currency. However, even if we accept that non-custodial wallets are essential to digital currency, the idea that offline payments are necessary for digital currency is certainly false. Users have an interest to possess and control digital assets directly, regardless of the context in which they spend them. Holding digital assets directly takes custodians out of outbound transactions, allowing users to avoid the risk that the custodians might extract fees, block certain (or all) transactions from taking place, or build profiles of the behaviour of their customers by monitoring the details of their transactions, such as time, location, and counterparty. Bearer wallets can be used in a variety of contexts, including but not limited to in-person POS transactions. Importantly, the benefits of non-custodial wallets can also be realised in a fully online context such as e-commerce.
Proponents of digital currency are right to acknowledge the secular decline in the use of cash in retail consumer transactions as a primary motivator for digital currency. As fewer consumers use cash for transactions, the variable revenues associated with operating a cash infrastructure are falling below the fixed costs. This pattern is evident in much of the developed world. For example, in the UK, banks are closing ATMs and branches [12], making cash more difficult for consumers to withdraw [13]. Some merchants have decided that the marginal cost of the various expenses related to accepting cash payments, such as tills, employee time, cash accounts with banks, and the means of transporting cash, outweighs the benefit of providing goods and services to cash-only customers. As a result, the value of cash to consumers is reduced, since customers no longer have assurance that their cash will be accepted at the point of sale, creating a feedback loop.
And so, when proponents of digital currency say that digital currency should allow “cash-like” transactions for retail users of money, they are referencing the idea that digital currency can provide individuals with a payment option that provides the key affordances of cash, including direct ownership, privacy, and equal access without discrimination. Users of digital currency would hold direct obligations of the central bank, and, given the right regulations, they might have reason to believe that their digital currency would be accepted by retailers that require electronic transactions.
Digital currency can certainly address problems associated with the systematic decline in the set of available options for consumers to access cash and the increasing tendency for merchants to refuse payments in cash at the point of sale. Users can load non-custodial wallets with the help of a personal device with network access, as well as via facilities provided by banks or ATM networks. Consumers can use non-custodial wallets in-person, to interface with point of sale devices in the same manner that they use debit cards or other custodial payment mechanisms today. Such payments can be considered “offline” both in the sense that they are in-person and in the sense that that the non-custodial wallets themselves do not require Internet connections. We refer to such payment scenarios as single-offline, because only one of the counterparties to a transaction has access to third parties via a network or other means of connectivity. Single-offline payments include in-person EMV card payments via POS systems and are therefore constitute a significant share of the payments that have replaced cash payments in the retail payments sector over the past twenty years.
2.3 Hold offline, transact online
The assumption that transactions that are not between accounts must be offline is wrong. This assumption would imply that all payments that are not online between payment accounts must be offline between peers that are disconnected from real-time communication with any other party. In particular, the proposal by the European Central Bank (ECB) for the digital euro [14] holds that these two models for payment, online via accounts and dual-offline between peers, are comprehensively exhaustive, but this is simply false. It simply ignores the possibility that money can be held offline and transacted online. Specifically, it is possible that a consumer would hold digital currency tokens in a non-custodial wallet and transfer them to a merchant with online connectivity, in much the same way that a consumer might rely upon the online connectivity of a POS device offered by the merchant.
There is a clear case for using digital currency with non-custodial wallets as a substitute for electronic card payments or network platform transactions at the point of sale. What remains is the question of whether there exists a need to support payments between two parties such that neither of which is connected to a network. There is no evidence that consumers who cannot withdraw cash or merchants who refuse to accept cash are using alternatives to cash that do not make use of a network connection instead, and there is no evidence of a significant demand for one, either. As a means of facilitating a transaction between two parties in situations in which neither party has a networked device, cash is remarkably effective. Although it is possible to imagine narrowly construed use cases for new capabilities, such use cases do not justify system requirements.
It is worth asking whether governments or central banks are harbouring an intention to do away with cash entirely, or whether the presumption of such an intention is shaping the view of what CBDC should be. For example, some authorities have argued that a larger share of domestic transactions should be brought inside the scope of the formal economy [15, 16]. Legitimate or not, that goal is not inconsistent with prevailing mechanisms to roll out digital transactions in general and does not require CBDC. Moreover, few monetary authorities or government regulators have called for the abolition of cash, whereas many, including those of Switzerland, the United Kingdom, and the United States, have pledged to preserve access to cash for the foreseeable future. Notably, Sweden has rolled back its drive toward creating a cashless society to support cash infrastructure in a near-cashless economy, at great cost [17]. So the idea that CBDC is mostly a justification for phasing out cash, as has been speculated by some, is not corroborated by facts.
3 Certified hardware as a de facto custodian
Design approaches based on certified hardware have featured prominently in the recent debate concerning the deployment of large-scale digital currency systems in general and systems to support CBDC in particular. Certified hardware is a cornerstone of the so-called “trusted computing” paradigm, wherein a user carries a device designed to operate in a manner that respects the will of a third party in preference to the will of the user. The justification for such approaches is usually given in terms of the perceived need to facilitate offline payments or to facilitate the recovery of lost assets. In this section, we examine the characteristics of trusted computing within the context of its application to digital currency systems and its potential impact on the power relationships between the users of devices, the manufacturers of devices, and other powerful actors such as corporations and the state.
Baressi and Zatti explicitly assume the use of trusted hardware to draw transactions into the security envelope of a trusted custodian [5], implicitly drawing the hardware manufacturer into the role of trusted third party. In their model, the wallet would be a means of transacting with money held in custodial accounts that would be reconciled ex post.
3.1 Fair exchange
If we define a fair exchange as a transaction mechanism that ensures that transacting parties either each receive or each do not receive what it expects (and do not receive “any additional information about the other’s item”) [19], then it is not possible to have a fair exchange between two parties without the involvement of a third party; this is a foundational truth [20]. With the exchange of physical objects, possession of the object itself is sufficient to demonstrate the veracity of the exchange that has taken place. But with digital objects the situation is more complicated. It is possible to have fair exchange without connection to the Internet, provided that there are third parties accessible locally, perhaps via a local network, as might be possible with ad hoc networks.
Proponents of approaches requiring certified hardware have proposed embedding into end-user devices a tamper-resistant “secure element” that can be used to systematically restrict what the device can do, disclose the immutable identity of the device to a third party, or furnish information to third parties about what the device had observed or done. The secure element would be impenetrable to the user, and the user would not be able to modify its behaviour without rendering the device inoperable. In effect, such an element effectively acts as the third party by proxy, by embedding the will of the third party into the operation of a device carried by a user.
3.2 The perils of trusted hardware
In the case of digital currency, the secure element would presumably be used as an on-board authority that would prevent double-spending by maintaining a record of transactions made by the device and disallowing users from transacting tokens that have already been signed over to others. This design model in particular, and trusted hardware in general, introduces a variety of problems. We consider some of those problems here:
  1. 1.
    Trusted computing introduces significant security risks particular to digital currency. The risks of relying upon the security of a hardware manufacturer are too great for a token-based CBDC context. In particular, a system design that relies upon trusted hardware components is fragile by design. If the trusted hardware fails to function as advertised or is compromised by an attacker, then users could potentially transact the same tokens an arbitrary number of times without bound. The system-level costs of such an attack would be huge, and the incentives would be tremendously enticing for attackers. We note that modern devices, such as debit cards, that rely upon tamper-proof hardware for the exchange of value generally interact with accounts. If their security is compromised, then the value available to attackers is limited by rules and operational procedures implemented by custodians. Without custodial accounts, such backup mechanisms are not possible, and so, a successful attack on the hardware root of trust would imply that it would be necessary for all users to replace their devices. Thus, it is reasonable to conclude that it is preferable to rely upon a protocol to achieve the required level of security, as can be done when one of the parties has access to the network, rather than to expect that trusted hardware modules would guarantee that all transactions would take place within the security envelope of a trusted custodian. To assume that all hardware wallets must take the latter approach because some circumstances might require it is tantamount to overspecification at the expense of users who would readily hold assets offline but transact them offline. It is worth noting that when Chaum, Grothoff, and Möser argue that “any economically producible device that stores tokens with monetary value in an individual’s possession, and enables offline transactions… will be the target of successful forgery attacks…” [34], they are referring to wallets that perform dual-offline transactions, in contrast to wallets that hold tokens offline and transact them online.
  2. 2.
    Secure elements are unaccountable to their users. Because trusted hardware is impenetrable to users, users cannot know with certainty what their devices are doing. The inability to analyse the behaviour of a device undermines the trust that users can and should place in the device. In essence, the device serves a second master that is not the user. Users cannot trust what they cannot verify, so they have every reason to be suspicious of unchecked promises on the part of hardware manufacturers or other privileged guarantors [21]. For example, secure elements can furnish information that can be used to identify the user or the user’s previous transactions. By contrast, with open hardware, skilled members of the community can conduct audits and provide reports to the public. Of course, users with the means to do so can personally conduct the audits themselves or hire contractors to conduct the audits for them. However, it is not necessary for every user to actually run an audit; evidence suggests that similar benefits are achieved so long as all users can perform audits if they choose to do so [22].
  3. 3.
    Certified hardware discourages innovation. Users who cannot create their own devices cannot innovate. Instituting a certification requirement for hardware serves to ensure that only those with the time and resources to pursue certification will be able to produce working devices. Regulations that increase the cost of entry are known to decrease entrepreneurship as well as innovation by existing firms [23]. Ordinary users and hobbyists will not be able to develop their own devices, creating a barrier to public understanding the of the core technology, encumbering public debate and consideration of alternative designs [24, 25].
  4. 4.
    Certified hardware distorts the market by privileging its manufacturers. The requirement to use trusted hardware effectively locks the hardware manufacturer in to a privileged position in the market, wherein prospective manufacturers require certification by some authority (either directly by the government or by a delegated authority such as a trade organisation), and concomitant lobbying and regulation can restrict free entry, increasing market concentration [26]. Furthermore, even if all manufacturers could have certification at zero marginal cost, the fact that fabrication facilities carry high fixed costs means that the market is likely to be highly concentrated, with manufacturers capturing the surplus [27].
  5. 5.
    A certified device is tantamount to a relationship with a custodian. Hardware manufacturers become gatekeepers at best, and globally trusted third parties in the general case. But unlike trusted third parties on the network, users bind themselves to long-term relationships with personal devices that contain trusted hardware; the relationships are not much different between account-holders and custodians, wherein account-holders are subject to satisfying the requirements of custodians. The cost of switching bank accounts is generally only a matter of time and effort on the part of account-holders, whereas the cost of switching devices generally includes per-device manufacturing and retailing cost plus taxes and costs associated with physical acquisition and disposal of the devices. In addition, while it is possible to implement both technical protocols and regulations that rely upon locally-trusted third parties, economies of scale mean that hardware manufacturers are likely to have global reach. It is reasonable to conclude that relationships between users and their devices featuring specific trusted hardware will be sticky indeed.
For these reasons, we are unconvinced that trusted hardware solutions are appropriate as requirements for the design of tools that allow users to possess and control their own assets. At the same time, we would not suggest that trusted hardware should be forbidden. In some fraction of cases in which physical cash is not an option and network connections are not available, trusted hardware might be an alternative kind of third party for transacting parties to consider, although this possibility is not sufficient justification to institute a requirement. In all cases, both sender and recipient must willingly choose to trust the third party to consummate the transaction, and given the final three points in the last above, it is difficult to imagine that a user preference for a particular device manufacturer would be real. Consider, for example, that existing users of secure hardware in the form of electronic payment cards generally have little say in which payment network their issuing banks choose.
3.3 Standards and lobbying
It is plain to see why a hardware manufacturer would lobby in favour of regulations and standards that declare “offline transactions” to be an essential requirement for CBDC. But this argument is unsound: Cash is an excellent offline payment method, and to our knowledge, few if any central banks have suggested that the purpose of CBDC is to replace cash outright. The argument implicitly equivocates between a concept of “offline” that denotes transactions wherein a user is not required to carry a device with network connectivity and a concept of “offline” that involves two counterparties physically meeting in a location without third parties or network infrastructure of any kind nearby. The former scenario is the only one in which users commonly use certified hardware today. The idea that cash could suddenly become unavailable for use in the latter scenario is an overly simplistic justification for hardware manufacturers to tout their products.
We note that secure elements can be used for a variety of purposes, including the prevention of exfiltration of private data from a device. However, requiring the use of secure elements for the purpose of creating attestations to parties other than the user of the device (i.e. “remote attestations”) that a device is operating within the security envelope of a trusted third party, whether a hardware manufacturer, service provider, or authority, implies a specific model for securing transactions that is not necessary for single-offline payments. For the reasons described above, we argue that single-offline payments would be both more robust and more secure if counterparties can trust network service providers, and the various methods network service providers can use to demonstrate the veracity of their claims, rather than having no choice but to rely upon trusted hardware physically held by the other counterparty. And so, for single-offline payments, trusted hardware adds nothing, but imposes significant costs, and therefore it must not be considered a requirement. Dual-offline transactions can certainly make use of remote attestation if both parties agree, but such transactions should not be seen as a general case for digital currency transactions, particularly given pervasive connectivity and the security benefits of using protocol-based security rather than introducing a requirement for counterparties to trust each other’s devices. For this reason, the development of standards that recommend or require hardware wallets to use a secure element for remote attestation are certainly problematic.
Standards that recommend hardware wallets to use a secure element to prevent the exfiltration of private data might be appropriate in the case of transactions involving smartphones, because of their complexity and large attack surface. However, it is inappropriate to assume that all or most digital wallets will be smartphones, not least because of their intrinsic security and privacy risks, despite being the “benchmark in user-accessible features”, as Barresi and Zatti argue [5].
4 Digital currency means tokens, not accounts
Barresi and Zatti assume that wallets would represent end-user money as account balances [5]. However, for ordinary consumers and non-financial businesses to hold digital currency directly, there must be some way for digital currency to be held outside accounts. Accounts generally involve a relationship between a beneficial owner (the account holder and a fiduciary, typically a regulated financial institution. A salient property of accounts is that access to assets held within accounts is determined by the identity of the account holder [28]. In contrast, tokens are assets that can be held outside of accounts. Tokens can be bearer instruments, meaning that they can be used in a way that does not index the identity of their owners. Payments can be held offline and transacted online, so the user experience would be similar to an EMV card payment except in which the payer presents digital assets rather than instructions to debit an account.
Compared to assets held within accounts, assets that are held directly by their owners have a different set of security requirements for their owners. For example, owners of assets held within accounts can rely upon various identification mechanisms, such as PIN numbers and two-factor authentication, to protect their assets if their cards or mobile devices are stolen, and asset custodians can also provide credit to account holders, such that if assets are spent unlawfully, the account holders can ask the asset custodians cover the value of their losses. In contrast, the bearer is responsible for bearer instruments. If cash is lost or stolen, owners generally have little recourse beyond law enforcement mechanisms to recover their money, and such mechanisms often fail to recover the value. The potential for loss or theft exists not only physically but also electronically. Devices can malfunction, and devices can be compromised. The prevailing model for handling devices that no longer work or have been compromised is to wipe their contents completely or to replace them wholesale. When such devices are used to access accounts, such replacement is not much of a problem, although bearer instruments inexorably expose their owners to new risks that require different solutions.
Some central banks, including the ECB [14], the US Federal Reserve [29], and the Bank of England [31], have presented arguments that holding limits for wallets are necessary for financial stability and to prevent bank runs. Such arguments are nonsense. Limiting CBDC withdrawals, rather than aggregate holdings, would be no less effective in preventing bank runs, and the approach of limiting bank withdrawals has been used successfully in the past, notably in Greece in 2015 [30]. Importantly, limiting CBDC withdrawals can be done without requiring wallets to implement account-like control mechanisms that undermine the will of their users. Barresi and Zatti suggest that wallets could perform “behavioural analytics” on their users [5]; little could be more disrespectful of a user’s autonomy.
Although a third party is necessary for fair exchange (and indeed trusted hardware or closed-source software blobs can be this third party de facto), Pagnia and Gärtner have shown that the role of the third party can be limited, even zero-knowledge: tokens can be transferred without the third party knowing what is being transferred or having a custodial relationship with either party [20].
The idea that digitalisation implies the use of accounts is not true in general. Digital tokens can be real indeed, with ownership demonstrable by transferring it: using a single-use private key to transfer the asset to another single-use private key. The requirement of an account in the form of a “wallet provider” is not true for cryptocurrencies in general, nor does a private key constitute a proof of identity when it is created to be used exactly once. It is true that many Bitcoin users rely on service providers to manage their holdings. However, to argue from this that self-custody of tokens will never take off is pure speculation, and it ignores the fact that most Bitcoin users are speculators with no expectation that they would use Bitcoin for ordinary retail purposes. Many people carry around an amount of cash that is useful for retail purchases but small enough that they could lose it without too much regret, despite the fact that losing it is forever.
Digital assets can certainly be held by users without accounts of any kind, for example, as a payer can in Chaum’s model. Accounts for recipients might be desirable, but they are not technically necessary. The involvement of a third-party to prevent double-spending does not imply that the third party must have a role as custodian or account provider, or knowledge of the asset being transferred, the circumstances of the transfer, or the identity of either party.
Tokens are also important for privacy. Systems that make use of transactions involving balances, even if the transactions themselves are privacy-preserving [36, 37], intrinsically require end-users to retain data produced during a transaction that can be used to prove that the transaction was carried out according to the rules. This means that transactions are not independent of one another. The property of transaction independence is important because retaining data related to past transactions introduces the risk that the user could be de-anonymised via poor operational security or blackmailed to reveal information about those transactions [38].
Finally, the idea that digital currency intrinsically requires accounts is not true; central banks can certainly issue tokens that exist outside of accounts, just as they do with cash, and a token like this would be a liability on a central bank’s balance sheet, not unlike cash or central bank reserves.
5 Separating issuance from transaction processing
Even if transactions are done with network access to a mutually trusted third party, there is no reason to assume that the issuer must be involved in transaction processing in real-time. Chaum’s original proposal assumed such a dependency [33], but this is not a requirement in general. In particular, it is possible to register a (zero-knowledge) commitment to transfer an asset without communicating with the issuer, and for a transaction to have finality it is sufficient for both parties have faith in a mutually trusted third party to eventually post the commitment to a bulletin board recognised by the issuer [10]. Also, if tokens carry their own provenance information, then there is no need for the issuer to track tokens that it had seen, either [10].
It is certainly possible to transact money issued by a central bank in a decentralised system, and wallets can be expected to respect this possibility. Trusting a central bank (or other issuing authority, such as those that issue e-money) to issue money and maintain its integrity is certainly different from trusting it to process transactions. In particular, it is possible to use a decentralised system to transact centrally issued money. The use of a secure element to establish trust implicitly requires the transacting parties to trust a hardware manufacturer or other central authority to facilitate transactions. Such a role for a unified central authority might be incompatible with the concept of a decentralised system. Finally, requirements for hardware wallets used for digital currency should not be needlessly inconsistent with requirements for hardware wallets used for other digital assets.
It is also worth considering that all users cannot be expected to trust the same trust anchors equally. Although the fact that assets held in offline non-custodial wallets can be transacted online is sufficient to reject the requirement of a hardware root of trust for security, even the case of dual-offline transactions poses a problem for requiring the use of secure elements. Specifically, there is the matter of who the transacting parties choose to trust. We imagine that parties choosing to transact online could potentially choose from a variety of manufacturers of hardware roots of trust based upon their needs and security models, whereas forcing manufacturers of hardware wallets to embed hardware certified by a specific authority, with the expectation that this authority is universally and equally trusted by all users, is not appropriate. We believe that allowing users to retain a choice of mutually trusted hardware security solutions, such as removable eSIM devices, is appropriate, although it might be impractical for users to carry around as many hardware devices as there are commonly trusted third parties with their various counterparties. The fact that hardware devices impose a trust model based upon their specific manufacturer is yet another reason to reject the idea that such devices could function as a public utility.
6 Wallets as containers for bearer instruments
For a set of specific technical requirements for digital currency in general to be credible, it must address the requirements of its various stakeholders. One of the unfortunate characteristics of the current debate is the diversity of conflicting assumptions about the nature of payments and the relative importance of different kinds of payment scenarios. In reality, there are many different ways that people make payments today: with and without custodians, locally and at a distance, with and without electronic devices, with and without money issued by a central bank. Money is transferred to pay invoices, to purchase goods and services, to give gifts, to pay taxes, or to reallocate funds within an organisation or family. These different scenarios imply a plethora of different costs and risk profiles: Buying a cup of coffee is different from buying a house. We should accept that there will not be a single payment system or key mechanism to rule them all, although it might be possible to develop a common framework that can accommodate a variety of different payment systems.
The case for trusted hardware, as for other core mechanisms of proposed payment systems, involves the requirement for assessing the integrity of money in a fair exchange scenario wherein counterparties lack access to a mutually trusted third party. In principle, trusted hardware can act as a trusted third party to enforce system-level rules on user devices directly, by implementing rules that prevent assets from being spent more than once. However, considering the security risks associated with compromised devices, it is better to enforce such rules over the network instead, if it is possible to do so. This can be done, for example, by invoking a trusted third party to validate all transactions, as proposed by Chaum and others [33, 34]. However, there is reason to believe the costs, risks, and public accountability concerns associated with having a central operator dictate that history unchecked are too great for the CBDC use case [35], and the most important costs and risks would be borne not only by the users but also by the central operator. A modern alternative involves using a distributed ledger system to put the money itself on a blockchain, the way that UTXO designs (e.g. Bitcoin [32], Monero [39]) and state-transition designs (e.g. Ethereum [40]) do. Such approaches are unwieldy because they require the ledger to track all of the tokens individually, and transaction finality requires global consensus. In practice, most cryptocurrency users rely upon custodial accounts, negating a foundational benefit of digital currency.
However, distributed ledgers and blockchains can be used to support and enable payment systems in a variety of ways, including by providing non-repudiable evidence that a transaction had taken place, even if an individual service provider might equivocate [41]. The digital currency design proposed by Goodell et al [10] and further elaborated by Friolo et al [38] addresses the scalability problems associated with modern cryptocurrency systems while strongly preventing double spending without trusted hardware. In particular, it avoids requiring the issuer to participate in every transaction of its tokens, it avoids requiring system operators to manage a list of tokens that have been created or spent, and it avoids requiring users to have access to a global network at the time of the transaction. Any third party chosen by the spender and accessible to the recipient at the time of the transaction can suffice.
Proposed solutions that rely upon stateful assets and oblivious transfers can work technically without distributed ledgers, but in the end, digital currency systems require a means to support an immutable, shared history mutually accepted among independent actors, which is the salient characteristic of a distributed ledger [42]. The immutability of a distributed ledger is emergent property of the independence of the participants. So, although using a distributed ledger to record and process tokens directly on the ledger might bring unacceptable costs in terms of scale, efficiency, and operational burden, we should not rule out using distributed ledgers for what they do best: externalising commitments so that parties responsible for maintaining the history do not equivocate [10]. These commitments can be aggregated in such a way that the ledger grows at a fixed rate determined in advance, and each transaction carries zero marginal cost to the ledger operators.
Trusted computing is not necessary to protect users from loss or theft, either. To protect their assets, users may, at their option, choose devices that offer security features such as encryption, passcodes, or local biometrics to protect against theft, and they may make copies of their tokens to mitigate the risk of accidental loss of their devices. Users can retain copies of their tokens, and they can share copies of the tokens with the police if the devices containing them are lost, stolen, or compromised. Users can also upload copies of their tokens to non-custodial service providers, such as cloud storage services.
A much better way to address the limited set of cases in which “fully offline” payments (that is, those where neither the payer nor the payee have access to any suitable third parties or network infrastructure) are necessary would be to support the use of cash itself where cash continues to work well, such as in remote locations and environments with low connectivity, while avoiding all of the aforementioned problems associated with trusted hardware. To address the reduction of cash in the “built” environments by providing digital currency, whilst encouraging the provision of facilities for converting between cash and diital currency at the boundary of “unbuilt” environments, would be an entirely legitimate approach. In particular, if one potential benefit of CBDC in an age of deteriorating cash infrastructure is the ability to electronically transport cash to remote population centres during crises [43], then there is no reason to require that central bank money distributed in this manner must remain digital. For example, it can be exchanged for physical tokens specific to that population centre, much as cash and bank money are exchanged for chips at a casino.
None of this is to say that we should forbid limited groups of users from trusting certified devices, if that is what they freely choose, to conduct transactions without network connectivity to a third party. Such an approach is entirely compatible with an architecture that expects that the third party would generally operate as a network service. However, at a system level, requiring all users to purchase and carry certified devices, even if some users intend only to hold money offline so that they can transact it online, introduces costs and risks that are entirely avoidable.
7 Conclusion
In the end, we must ask ourselves: What problem are we really trying to solve with digital currency? Manifestly, the problem is the decline of cash, and the concomitant challenges of maintaining cash infrastructure when the variable revenues of cash infrastructure fall below the fixed costs. But the decline of cash really only exists in environments with high network connectivity, wherein dual-offline payments are not necessary. For environments with low network connectivity, the decline of cash has not occurred, and cash continues to be the best solution for such environments.
Many central banks, including the Bank of England, have affirmed a commitment to cash, while underscoring the risks associated with the decline of cash use in favour of online payment schemes such as card payment networks. Really, the use cases that digital currency payments should be targeting are cases for which at least one counterparty is already online. To argue for digital currencies with the premise that support for dual-offline payments are its principal source of added value mainly supports the view that such products are solutions in search of a problem.
The purpose of digital wallets (and indeed, digital currency in general) is not primarily about enabling transactions in which both parties are offline. The purpose of digital wallets is to allow users to hold assets directly, without a custodian. Put differently, this is a way for users to have private property that they can exchange. By facilitating such payments, digital currency can protect the interests of consumers in all cases, including fully networked ones. If the counterparty has network connectivity (for example, an electronic point of sale device), then transactions involving digital wallets can be consummated without special hardware. Even if we accept that fair exchange is impossible without the participation of a trusted third party, and that the trusted third party can participate via the network or via trusted hardware, it is important to underscore that a non-networked digital wallet can conduct a fair exchange with a networked counterparty, without the aid of special security mechanisms built into the wallet.
It is also a mistake to assume the primary motivation for CBDC is to allow a government to phase out cash, or that custodial wallets are sufficient for all scenarios except the narrow case in which neither the sender nor the receiver have a communication line to a third party. Part of the reason that cash is successful is its role as a bearer instrument. Credit systems actually predated cash, and cash emerged as a partial substitute [44]. What the world does not yet have is an institutionally supported version of cash for the digital economy, and a digital currency aspiring to fill this role will require a non-custodial wallet architecture.
Custodial wallets require network access as a rule, because consumers must be able to use their devices to access their custodial accounts. However, digital wallets do not imply a design choice between network access and certified hardware in the general case. In particular, single-offline payments can make use of non-custodial wallets without certified hardware to hold money offline and transact it online, leveraging the network connectivity provided by POS devices presented to customers by merchants. To deny the salience of this use of wallets is to impede the development of a technology that may one day become crucial for the digital economy. Ultimately, the digital wallets that support economic transactions should be seen not as access devices or proxies offered by custodians, but as tools freely chosen, and even built, by end-users to store and use digital money autonomously, on their own terms.
Acknowledgements
The author thanks Professor Tomaso Aste for his continued support for our work on digital currencies and digital payment systems. The author also thanks his many colleagues in the international standards development community for a series of vigorous arguments that were instrumental to reaching the crux of this most contentious debate. The author also acknowledge the Future of Money Initiative at University College London, the Systemic Risk Centre at the London School of Economics, EPSRC, and the PETRAS Research Centre EP/S035362/1 for the FIRE Project.

### v4

@@
 In this article, we critically assess the essential problems that digital currency solutions are being proposed to solve, particularly with respect to the future of payments and the future of cash.
-We assess the validity of common justifications for account-based payments and certified hardware in the context of alternative designs, limitations, and trade-offs.
-We conclude that the interests of consumers would be better served by design approaches to digital currency that anticipate that digital assets would be held outside accounts, stored offline, but transacted online, without requiring the use of trusted hardware.
+We evaluate the validity of common justifications for account-based payments and certified hardware in the context of alternative designs, limitations, and trade-offs.
+We conclude, referencing some specific designs, that the interests of consumers would be better served by design approaches to digital currency that anticipate that digital assets would be held outside accounts, stored offline, but transacted online, without requiring the use of trusted hardware.
 1 Introduction
-The proliferation of cashless payments has given rise to a growing preponderance of merchants who refuse to accept cash, as well as the suggestion that private-sector payment systems could subvert the authority of central banks to implement monetary policy.
+The proliferation of cashless payments has given rise to a growing number of merchants who refuse to accept cash, a well-documented and legal practice in some jurisdictions around the world [De Nederlandsche Bank 2023, Grabar 2018] that has correspondingly given rise to government reports and consultations [House of Commons 2025, Australia 2024].
+Combined with the emergence of novel private-sector payment systems, the decline of cash subverts the authority of central banks to implement monetary policy.
 This development has, in turn, focussed the attention of central banks and other financial institutions around the world on digital currency, particularly the idea that central banks could issue central bank digital currency (CBDC) for retail use.
 In principle, retail CBDC would be a direct obligation of the central bank, like cash or central bank reserves, but existing in digital form and held by ordinary consumers and non-financial businesses.
-Many proponents of retail CBDC have argued that it must be “cash-like” to deliver value to end-users.
+Some proponents of retail CBDC, including central banks, have argued that it should have cash-like properties to deliver value to end-users [ECB 2021, Bank of England 2023].
 This argument is at least partially right.
-For example, it might be expected that a cash-like system would afford users privacy by design, resistance to discrimination and profiling, and the ability to possess and control the assets that they use to engage with the economy.
+For example, it might be expected that a cash-like system would afford users privacy by design, resistance to discrimination and profiling, and the ability to possess and control the monetary assets that they use to participate in the economy and satisfy their economic needs.
 These are certainly desirable properties.
-However, some authorities have interpreted the argument that a digital payment system must be cash-like to mean that a digital payment system must enable dual-offline transactions, namely those without connectivity of any kind between either of the counterparties and any third parties, in contrast to “single-offline” transactions, wherein a mutually trusted third party is reachable by one of the counterparties.
-We suggest that this argument constitutes a leap of logic that must be examined carefully, starting with its implicit assumptions that any new payment system must be appropriate for all scenarios, that existing solutions such as cash are insufficient to handle payment scenarios in which neither party has access to a network connection, and that there is a specific need for future digital currency infrastructure to enable a kind of digital transaction that is not commonly available to users today.
-Importantly, the idea that digital currency transactions must mirror the current system for electronic retail payments demonstrates an assumption that the main promise of digital currency concerns accessing custodial accounts whilst offline.
-But this assumption ignores decades of history of digital currency, wherein the most groundbreaking systems, including not only Bitcoin but also DigiCash two decades earlier, drew attention by introducing a way to hold money outside of custodial accounts.
+However, some authorities have interpreted the argument that a digital payment system must be cash-like to mean that a digital payment system must enable “dual-offline” transactions, namely those without connectivity of any kind between either of the transacting parties and any third parties, in contrast to “single-offline” transactions, wherein a mutually trusted third party is reachable by one of the counterparties.
+The “dual-offline” argument constitutes a leap of logic that must be examined carefully, starting with its implicit assumptions that any new payment system must be appropriate for all scenarios, that existing solutions such as cash are insufficient to handle payment scenarios in which neither party has access to a network connection, and that there is a specific need for future digital currency infrastructure to enable a kind of digital transaction that is not commonly available to users today.
+Importantly, the idea that digital currency transactions must mirror the current system for electronic retail payments demonstrates an assumption that all digital money must take the form of an account offered by some service provider who acts as a custodian, and that the main promise of digital currency concerns accessing such an account for use in dual-offline payments (an “offline account”).
+Following this logic, a digital currency hardware wallet would be a tool by which an authorised account holder could conduct peer-to-peer transactions between devices in the absence of network connections.
+To achieve this, the wallet itself must enforce a set of rules, for example, the wallet must track the size of the account balance, and the amount of money that the account holder can spend must be limited to that value.
+However, there is another, competing view that maintains that the purpose of digital currency is to provide a digital form of cash (“e-cash”), and this view is supported by past digital currency design proposals over the past four decades.
+Some of the most groundbreaking systems, including not only Bitcoin [Nakamoto 2009] but also DigiCash [Pitta 1999] two decades earlier, drew attention by introducing a way to hold money outside of custodial accounts.
+Both of these systems were sometimes described as “cryptocurrency” because they were engineered using cryptographic techniques.
+Importantly, the tokens that users held were explicitly designed to be cash-like, as can be inferred from their names, and intended to be used for payments and not linked to accounts.
+In the case of Bitcoin, tokens were created and spent using an immutable ledger (in the form of a blockchain system), and in the case of DigiCash, tokens were created and spent as bearer instruments using a redemption process with a central operator, which is assumed to be responsible and accountable for its issuances, which can be held anonymously by prospective payers.
 The idea that central banks would never support such an architecture ignores the fact that central banks have always supported the issuance and circulation of tokens outside accounts, in the form of cash.
 Within this framework, offering cash in digital form is the next logical step.
-Dual-offline transactions impose additional security requirements to prevent double-spending attacks, in addition to the new security requirements applicable to digital bearer instruments in general.
-Some prominent experts have proposed “trusted computing” [1, 2] as a solution to the security problem, reasoning that if the security of bearer instruments is a cause for concern, then perhaps there is benefit to be realised in making an effort to ensure that the devices that can hold them must meet a certain standard.
+We suggest that the “e-cash” vision of digital currency is largely absent from the discussion of the design of digital currency hardware wallets, and we argue that this neglect is a mistake.
+Contemporary examples of “e-cash” systems with significant deployments include GNU Taler [GNUnet and Taler] and (de facto) Privacy Pass [Davidson 2018], both of which feature the use of non-custodial wallets to hold tokens for redemption.
+In both the “offline account” and “e-cash” use cases, dual-offline transactions impose additional security requirements on the transacting parties, in addition to the security requirements applicable to digital transactions in general.
+In particular, payers must be prevented from falsely claiming that the money they are using to pay has not previously been spent (a “double-spending attack”).
+Some prominent experts have proposed “trusted computing” [Trusted Computing Group 2004, Roberts 2021] as a solution to the security problem, reasoning that if the possibility that a payer (or a payer’s device) might misbehave, then perhaps there is benefit to be realised in making an effort to ensure that the devices that can hold them must be able to enforce rules directly, a “hardware root of trust”.
 After all, many users are accustomed to carrying certified hardware, and the prevalence of certified hardware among common user devices might be seen as a justification for their use in the context of digital currency.
-For example, smart cards commonly rely upon tamper-resistant chips containing secrets issued by their manufacturers [3].
+For example, smart cards commonly rely upon tamper-resistant chips containing secrets issued by their manufacturers [ETSI 2019].
 Mobile phones often contain secure enclaves to restrict what their users can do with their devices, for example to prevent users from copying multimedia files.
+E-cash systems are designed to work without tracking the ownership of individual tokens on a ledger.
+A generic asset “template” containing information such as a public key is first created by a prospective consumer, who then uses a zero-knowledge protocol to request a privacy-preserving certification of the digital asset, such as a blind signature, from the issuer, which then becomes part of the asset.
+The issuer can now recognise the certified digital asset as valid without recognising its owner or the circumstances of its creation.
+The asset is stored by the end-user directly and is not recorded on a ledger.
+Double-spending can be prevented in a few ways, one of which is to assume (a) that the issuer processes all transactions as redemptions, (b) that the issuer maintains a ledger to keep track of the set of redeemed assets that it receives in these transactions, and (c) that the issuer, upon receiving an asset in a transaction, checks the ledger to verify that the asset has not already been redeemed [Chaum 1982, Chaum 2021].
+An alternative approach, which we call oblivious management, does not require the issuer to process transactions or maintain a ledger of redeemed tokens.
+In such a system, the assets carry their own provenance information, linking transaction information they carry to rolled-up aggregation entries, one per time period, on a ledger [Goodell 2023, Friolo 2025]; however, neither the tokens nor the value they represent are represented on the ledger at all.
+In both cases, issuers might keep accounting records of the tokens that they issue or redeem, but the privacy-preserving certification unlinks the issuance of an asset, where the party requesting a new asset may be identified, from its subsequent payment transaction and redemption, wherein only the payee is known.
 We also consider the conflation of technology, which can be freely and directly deployed by an individual or business to perform a function, with services, which are functions performed by third parties.
 In a world of smartphone applications for which most of the computation is nearly always performed by remote service providers, it is easy to ignore this distinction.
-However, since the ability to possess and control money directly is essential to its function, we must be careful to distinguish between technology that is under the control of the end-user and technology that is under the control of a third party.
-Of course, the fact that certain methods are commonly used in some contexts does not justify their use in others, and it certainly does not justify mandating the use of such methods in the context of future infrastructure that intrinsically must be available for general use by the public.
+However, since the ability to possess and control cash directly is functional to its use, we must be careful to distinguish between technology whose functions are governed by the end-user and technology whose functions are governed by a third party.
+Of course, the fact that certified hardware controlled by third parties is commonly used in some contexts does not justify its use in others.
+It certainly does not justify mandating the use of such methods in the design of future public infrastructure, especially if doing so would compel users to relinquish some of their control to third parties.
 It is all too easy for those in the payments industry to assume that everyone has or should have a smartphone and a bank account.
-Mandating the use of certified hardware carries many costs, and the benefits of certified hardware are not as clear as its proponents claim.
+Mandating the use of certified hardware carries many costs, and the benefits of certified hardware are not as clear as its proponents claim [Giesecke+Devrient 2023, Thales 2020].
+In particular, we suggest that much of the conversation in Europe about digital currency wallets has followed from the expectation, perhaps inspired by the digital euro regulation proposed by the European Commission [European Commission 2023], that for an institutionally accepted digital currency in general, and central bank digital currency in particular, the salient feature of cash is that it can be transferred in a peer-to-peer way.
+We disagree.
+We argue that the salient feature of cash is that it is a bearer instrument, and because peer-to-peer transactions are not a necessary characteristic of bearer instruments, we argue that it is worth considering protocols that do not rely upon wallets to enforce the rules, but for which the locus of enforcement lies with service providers.
+This argument has separately been articulated by the GNU Taler project [Grothoff and Dold 2021].
+Even if such protocols rely upon at least one party to be online, they can be considered social-welfare enhancing with respect to protocols that rely upon users holding devices that they do not control, particularly given the well-established vulnerabilities in trusted hardware [GNU Taler 2021], and notwithstanding assertions offered by the ECB that some hardware technology, such as embedded secure elements and SIM devices, are more robust against attacks than trusted execution environments [ECB 2025].
 The aforementioned considerations raise the question of how end-users will hold digital currency assets, and in particular, what the role of a wallet should be.
-ISO 22739:2024 defines a wallet as an “application or mechanism used to generate, manage, store or use private keys and public keys or other digital assets” [42].
-The simplicity of this definition would seem to understate the potentially complex features required to implement token management and custody, an exploration of which is given by Barresi and Zatti [5].
+ISO 22739:2024 defines a wallet as an “application or mechanism used to generate, manage, store or use private keys and public keys or other digital assets” [ISO 22739 2024a].
+The simplicity of this definition would seem to understate the potentially complex features required to implement token management and custody, an exploration of which is given by Barresi and Zatti [Barresi 2023].
 We address some of the assumptions that Barresi and Zatti make about the design of wallets and suggest different ways of viewing the problem.
-In particular, we critically address the arguments that trusted computing is an indispensable enabler of digital currency solutions, that digital currency must be held in accounts, that wallets must be linked to the identity of a user, and that a wallet is properly seen as “not a place where you can store cash [but] that helps you keep track of it” [5].
+In particular, we critically address the arguments that dual-offline payments constitute the most important use case for digital currency (see Section 2), that trusted computing is an indispensable enabler of digital currency solutions (see Section 3), that digital currency must be held in accounts (see Section 4), and that a wallet is properly seen as “not a place where you can store cash [but] that helps you keep track of it” [Barresi 2023] (see Section 6).
+Additionally, we address the argument implicit to some e-cash solutions that issuance and transaction processing must be performed by the same actor (see Section 5).
+Throughout this article, we shall reference a set of specific designs, including e-cash designs, such as GNU Taler, that are based on bearer instruments, as well as a proposal for how to integrate bearer instruments into a mechanism for using an integrity provider to transfer digital assets obliviously, without maintaining records of the assets, without immediately redeeming them for the value they represent, and without recording any information about the assets or the counterparties involved in transactions [Goodell 2023, Friolo 2025].
 The rest of the article is organised as follows.
@@
 2 The sophistry of dual-offline payments
-Recently, several prominent organisations, including the International Monetary Fund [6] as well as central banks such as the Bank of England [7] and the Bank of Canada [8], have released reports suggesting that dual-offline payments are desirable or even essential to the functionality of a future retail CBDC system.
+In addition to the ECB [ECB 2025], several prominent organisations, including the International Monetary Fund [Kiff 2022] as well as other central banks, such as the Bank of England [Bank of England 2022] and the Bank of Canada [2], have released reports suggesting that dual-offline payments are desirable or even essential to the functionality of a future retail CBDC system.
 Such bold statements are striking, considering the prevailing lack of consensus about what the requirements for digital currency should be in general.
-In this section, we shall critically explore the argument that the purpose of digital currency is to facilitate access to custodial accounts in a dual-offline payment scenario.
+In this section, we shall critically explore the argument (or implicit assumption) that the purpose of digital currency is to facilitate access to custodial accounts in a dual-offline payment scenario.
 2.1 Kinds of payments
@@
 They represent a significant and growing share of the total value consumers spend in the economy, and naturally, they require consumers to be online.
+(An example is a consumer paying for goods or services by inputting card details into a form on a web page.)
   • Bank transfers.
@@
 Settlement between the banks is managed via an interbank clearing and settlement network.
-In most cases, interbank settlement networks cannot guarantee “immediate” settlement within a few seconds, and bank transfers are usually not considered appropriate at the point of sale or for transactions that require acceptance in real-time.
+In some cases, interbank settlement networks cannot guarantee “immediate” settlement within a few seconds, and, despite the emergence of new open banking architectures, account-to-account “push” payments (in contrast to “pull” payments, such as via card networks) have yet to see global adoption at the point of sale or for transactions that require acceptance in real-time.
+(An example is a consumer instructing her bank to pay an invoice, either over the phone, in a branch, or with a banking app or website.)
   • Electronic card payments at the point of sale (POS).
 Electronic card payments provide a mechanism by which merchants can accept in-person payments from consumers in real-time.
-The consumer’s bank (the issuer) and the merchant’s bank (the acquirer) both participate in a payment network.
-The issuer provides a card to the consumer, and the acquirer offers to connect a merchant’s POS device to the payment network.
+The consumer’s bank (the issuer) and the merchant’s bank (the card payment acquirer) both participate in a payment network.
+The issuer provides a card to the consumer, and the acquirer provides a POS device to connect the merchant to the payment network.
 The card serves as a form of identification: Its purpose is to specify a particular bank account and to identify its user as the beneficiary of that account.
 To protect against unauthorised copying of the identification credential, cards are usually designed with features that provide tamper-resistance.
-The client is not assumed to have an Internet connection, and the POS device provides a way for the card to authenticate itself to the payment network.
-If the authentication is successful, then the payment network effectively underwrites the transaction, implicitly providing credit to the issuer and prompting the acquirer to accept the transaction immediately.
+The client is not assumed to have a real-time connection to an acquiring bank or service provider, and with contemporary ISO 8583 payments [ISO 8583 2023], the POS device provides a way for the card to authenticate itself to the payment network.
+If the authentication is successful, then the payment not immediately settled, and an inter-bank credit is implicit when the acquirer accepts the transaction.
+When the transaction is accepted, the acquirer has a credit against the issuer, which owes the transaction amount to the acquirer.
+In the case of a debit card transaction, the issuing bank customer is debited immediately; in the case of a credit card transaction, the issuing bank customer gets a credit in exchange for incurring a debt against the issuing bank (credit card debt).
 The payment network informs the merchant that the transaction is approved, and the merchant can then accept the payment in real-time.
 Merchants pay interchange fees to support this mechanism.
+(An example is a consumer buying a cup of tea by presenting a card to a merchant POS device.)
   • Closed-loop payment systems at the point of sale.
@@
 Crucially, banks are not generally involved in transactions; the platform operator manages the process.
+(An example is a consumer using an Alipay or WeChat app to purchase goods or services using e-money credits.)
   • Cash.
@@
   • Can a transaction to be provisionally allocated using the network but consummated when both parties are disconnected from the network?
-Notwithstanding some measure of understandable confusion and inconsistency in the use of the term “offline” accompanying calls for considering offline payments to be an essential feature of digital currency, it can be said that some measure of online technology supports virtually all retail payments today, excepting only payments in cash, which has seen a dramatic decline in use throughout the world over the past decade.
-With this in mind, it is reasonable to ask: What problem do proponents of retail CBDC intend to solve, for which dual-offline payments constitute a better solution than online payments involving assets held offline?
+Payment type payer online? payee online? bearer token? dual-offline?
+E-commerce yes yes no no
+Bank transfer yes no no no
+POS card payment no yes no no
+POS closed-loop payment maybe yes no no
+Cash no no yes yes
+E-token bearer payment maybe maybe yes no
+Table 1: Features of different payment types.
+Notwithstanding some measure of understandable confusion and inconsistency in the use of the term “offline” accompanying calls for considering offline payments to be an essential feature of digital currency, it can be said that some measure of online technology supports virtually all retail payments today, excepting only payments in cash, which has seen a decline in use throughout the world over the past decade, starting well before the pandemic of 2020–2022.
+Table 1 summarises the various types of payment, including the proposed “e-token bearer payment”, which is a digital online payment using a bearer instrument that is currently not an option for most consumers.
+It is a payment type like this that could be considered the next step in the evolution of cash, cash for the digital economy.
+Like cash, this new bearer instrument would be possessed and controlled directly by its users, as an assignable obligation of the issuer.
+The issuer, which could be a central bank, would be responsible for fulfilling this obligation, but it would not necessarily hold collateral corresponding to the total volume of its issuance.
+With this in mind, it is reasonable to ask: What problem do proponents of retail CBDC intend to solve, for which dual-offline payments constitute a better solution than online payments involving assets that are held offline?
 We acknowledge that there are some limited real-world use cases for dual-offline payments, such as payments aboard transportation systems that lack reliable network connectivity throughout the system [1].
 However, such examples are rare indeed, owing to the prevalence of network connectivity in locations suitable for digital transactions.
-Furthermore, we note that payment scenarios similar to this example do not actually require dual-offline payments; their requirements could also be satisfied by time-shifting, in which consumers purchase e-money tokens offline and consummate the transactions offline by selectively revealing proofs of purchase [10].
-And so, it is fair to ask: Might the idea of dual-offline payments constitute a solution in search of a problem, an ephithet that a report by the UK House of Lords prominently ascribed to CBDC in general [11]?
-2.2 Digital currency use cases
-One might argue that conducting transactions using digital currency held by custodians is, from the perspective of users and counterparties, not much different than conducting transactions practice of using ordinary bank accounts, with a possible exception of a different regulatory environment for the digital currency custodian.
+Furthermore, we note that payment scenarios similar to this example do not actually require dual-offline payments; their requirements could also be satisfied by time-shifting, in which a consumer purchases e-money tokens online and consummates transactions with a predetermined merchant offline, by selectively sending to the merchant tokens for which the consumer had already transferred control to the merchant [Goodell 2023].
+(The consumer and the merchant can use a fair exchange protocol to recover any unspent tokens later.) And so, it is fair to ask the question implicit to the title of a report by the UK House of Lords, “Central bank digital currencies: a solution in search of a problem?” [House of Lords 2021].
+2.2 Motivations for digital currency
+One might argue that conducting transactions using digital currency held by custodians is, from the perspective of users and counterparties, not much different than conducting transactions through ordinary bank accounts, with a possible exception of a different regulatory environment for the digital currency custodian.
 The custodian is still a gatekeeper that could fail to meet its obligations; the user is still bound to the custodian by identification requirements and still does not hold the assets directly.
-One might also argue that digital currency held by custodians is even less useful than bank deposits, if it is assumed that digital currency tokens cannot be rehypothecated and therefore cannot contribute to money creation.
+One might also argue that digital currency is less useful than cash if it is assumed that it can held only by depository institutions, or that individuals cannot deposit and withdraw it from depository institutions, or if it cannot be exchanged by depository institutions for central bank reserves.
 We might reasonably conclude that non-custodial wallets, which allow users to hold digital assets directly, are a foundational benefit of digital currency.
-However, even if we accept that non-custodial wallets are essential to digital currency, the idea that offline payments are necessary for digital currency is certainly false.
-Users have an interest to possess and control digital assets directly, regardless of the context in which they spend them.
+However, even if we accept that non-custodial wallets are essential to digital currency, the idea digital currency held offline must be spent offline is certainly false.
+Users have an interest to possess and control digital assets on devices of their choosing, rather than devices associated with attributable accounts, regardless of the context in which they spend them, for reasons of privacy and censorship resistance [Kahn 2005], which we shall address in Section 3.
 Holding digital assets directly takes custodians out of outbound transactions, allowing users to avoid the risk that the custodians might extract fees, block certain (or all) transactions from taking place, or build profiles of the behaviour of their customers by monitoring the details of their transactions, such as time, location, and counterparty.
 Bearer wallets can be used in a variety of contexts, including but not limited to in-person POS transactions.
-Importantly, the benefits of non-custodial wallets can also be realised in a fully online context such as e-commerce.
-Proponents of digital currency are right to acknowledge the secular decline in the use of cash in retail consumer transactions as a primary motivator for digital currency.
-As fewer consumers use cash for transactions, the variable revenues associated with operating a cash infrastructure are falling below the fixed costs.
-This pattern is evident in much of the developed world.
-For example, in the UK, banks are closing ATMs and branches [12], making cash more difficult for consumers to withdraw [13].
-Some merchants have decided that the marginal cost of the various expenses related to accepting cash payments, such as tills, employee time, cash accounts with banks, and the means of transporting cash, outweighs the benefit of providing goods and services to cash-only customers.
-As a result, the value of cash to consumers is reduced, since customers no longer have assurance that their cash will be accepted at the point of sale, creating a feedback loop.
-And so, when proponents of digital currency say that digital currency should allow “cash-like” transactions for retail users of money, they are referencing the idea that digital currency can provide individuals with a payment option that provides the key affordances of cash, including direct ownership, privacy, and equal access without discrimination.
-Users of digital currency would hold direct obligations of the central bank, and, given the right regulations, they might have reason to believe that their digital currency would be accepted by retailers that require electronic transactions.
-Digital currency can certainly address problems associated with the systematic decline in the set of available options for consumers to access cash and the increasing tendency for merchants to refuse payments in cash at the point of sale.
+Importantly, consumers can realise the privacy and anti-censorship benefits of non-custodial wallets in a fully online context such as e-commerce, for example, if they are using a system that supports asymmetric privacy [Tinn 2024].
+Proponents of CBDC are right to acknowledge the secular decline in the use of cash in retail consumer transactions as a primary motivator for digital currency.
+For example, in the UK, banks are closing ATMs and branches [Access to Cash Review 2019], making cash more difficult for consumers to withdraw and for businesses to manage cash [Tisher 2020].
+Although court judgments in some jurisdictions suggest that the legal tender status of banknotes implies mandatory acceptance in all but exceptional circumstances [European Union 2021], jurisprudence in other jurisdictions only recognises legal tender as something that must be accepted in the settlement of a debt [Bank of England 2025], admitting the possibility that vendors can freely choose what forms of payment they will accept in the general case.
+And so, when proponents of CBDC say that CBDC should allow “cash-like” transactions for retail users of money, they are foreshadowing the idea that CBDC can provide individuals with a payment option that provides the key affordances of cash, including direct control, privacy, and equal access without discrimination.
+Users of digital currency would hold direct obligations of the central bank, as they do with banknotes issued by the central bank, and, given the proper regulations, they might have reason to believe that their digital currency would be accepted by retailers that prefer electronic transactions.
+CBDC can certainly address problems associated with the systematic decline in the set of available options for consumers to access cash and the increasing tendency for merchants to avoid payments in cash at the point of sale.
 Users can load non-custodial wallets with the help of a personal device with network access, as well as via facilities provided by banks or ATM networks.
@@
 We refer to such payment scenarios as single-offline, because only one of the counterparties to a transaction has access to third parties via a network or other means of connectivity.
-Single-offline payments include in-person EMV card payments via POS systems and are therefore constitute a significant share of the payments that have replaced cash payments in the retail payments sector over the past twenty years.
+Single-offline payments include in-person EMV card payments via POS systems and therefore constitute a significant share of the payments that have replaced cash payments in the retail payments sector over the past twenty years.
 2.3 Hold offline, transact online
@@
 This assumption would imply that all payments that are not online between payment accounts must be offline between peers that are disconnected from real-time communication with any other party.
-In particular, the proposal by the European Central Bank (ECB) for the digital euro [14] holds that these two models for payment, online via accounts and dual-offline between peers, are comprehensively exhaustive, but this is simply false.
+In particular, by stating that “digital euro users should have the choice to use the digital euro either online or offline, or both”, the proposal by the European Central Bank (ECB) for the digital euro [European Commission 2023] implicitly suggests that the two models it describes for payment, online via accounts and dual-offline between peers, are comprehensively exhaustive, but this is simply false.
 It simply ignores the possibility that money can be held offline and transacted online.
 Specifically, it is possible that a consumer would hold digital currency tokens in a non-custodial wallet and transfer them to a merchant with online connectivity, in much the same way that a consumer might rely upon the online connectivity of a POS device offered by the merchant.
-There is a clear case for using digital currency with non-custodial wallets as a substitute for electronic card payments or network platform transactions at the point of sale.
+Given the rise of e-commerce transactions, the tendency of merchants to refuse cash at the point of sale, and the legitimate interest of consumers in privacy and censorship avoidance, there is a clear case for using CBDC with non-custodial wallets as a substitute for electronic card payments or network platform transactions at the point of sale.
 What remains is the question of whether there exists a need to support payments between two parties such that neither of which is connected to a network.
-There is no evidence that consumers who cannot withdraw cash or merchants who refuse to accept cash are using alternatives to cash that do not make use of a network connection instead, and there is no evidence of a significant demand for one, either.
+There is no evidence that consumers and merchants who dislike cash are seeking alternative payment technologies that do not make use of a network connection.
 As a means of facilitating a transaction between two parties in situations in which neither party has a networked device, cash is remarkably effective.
-Although it is possible to imagine narrowly construed use cases for new capabilities, such use cases do not justify system requirements.
+Although it is possible to imagine narrowly construed use cases for dual-offline payments, such as transient network outages, those use cases do not justify system-level requirements for a hardware root of trust on all end-user devices.
 It is worth asking whether governments or central banks are harbouring an intention to do away with cash entirely, or whether the presumption of such an intention is shaping the view of what CBDC should be.
-For example, some authorities have argued that a larger share of domestic transactions should be brought inside the scope of the formal economy [15, 16].
-Legitimate or not, that goal is not inconsistent with prevailing mechanisms to roll out digital transactions in general and does not require CBDC.
-Moreover, few monetary authorities or government regulators have called for the abolition of cash, whereas many, including those of Switzerland, the United Kingdom, and the United States, have pledged to preserve access to cash for the foreseeable future.
-Notably, Sweden has rolled back its drive toward creating a cashless society to support cash infrastructure in a near-cashless economy, at great cost [17].
-So the idea that CBDC is mostly a justification for phasing out cash, as has been speculated by some, is not corroborated by facts.
+The argument that a larger share of domestic transactions should be brought inside the scope of the formal economy is certainly not new [FATF 1996, Camera 2001].
+Recently, the governments of certain countries, including India [Ministry of Finance 2025] and Indonesia [Septiningrum 2024], have taken steps toward eliminating cash entirely, an objective that is consistent with prevailing business-led approaches to phase out cash transactions and encourage digital transactions in general.
+However, achieving that goal, whether it is legitimate or not, does not require CBDC; some countries have already implemented regulations to prevent the use of cash for large transactions or to control bank deposits comprising large amounts of cash [Financial Conduct Authority 2023].
+Moreover, few monetary authorities or government regulators have called for the abolition of cash, whereas many, including those of Switzerland [Reuters 2024], the United Kingdom [Bank of England 2024], and the United States [Federal Reserve 2022], have pledged to preserve access to cash for the foreseeable future.
+Notably, Sweden has rolled back its drive toward creating a cashless society to support cash infrastructure in a near-cashless economy, at great cost [Library of Congress].
+So the idea that CBDC is mostly a justification for phasing out cash, as has been speculated by some, is not corroborated by the facts, not yet at least.
 3 Certified hardware as a de facto custodian
 Design approaches based on certified hardware have featured prominently in the recent debate concerning the deployment of large-scale digital currency systems in general and systems to support CBDC in particular.
-Certified hardware is a cornerstone of the so-called “trusted computing” paradigm, wherein a user carries a device designed to operate in a manner that respects the will of a third party in preference to the will of the user.
-The justification for such approaches is usually given in terms of the perceived need to facilitate offline payments or to facilitate the recovery of lost assets.
+Certified hardware is a cornerstone of the so-called “trusted computing” paradigm, wherein a user carries a device designed to operate in a manner that respects the will of a third party rather than the will of the user.
+The justification for such approaches is usually given in terms of the perceived need to facilitate offline payments or to facilitate the recovery of lost digital assets (for instance, in case of loss or theft of the hardware device in which those assets were stored).
 In this section, we examine the characteristics of trusted computing within the context of its application to digital currency systems and its potential impact on the power relationships between the users of devices, the manufacturers of devices, and other powerful actors such as corporations and the state.
-Baressi and Zatti explicitly assume the use of trusted hardware to draw transactions into the security envelope of a trusted custodian [5], implicitly drawing the hardware manufacturer into the role of trusted third party.
+Barresi and Zatti explicitly assume the use of trusted hardware to draw transactions into the security envelope of a trusted custodian [Barresi 2023], implicitly drawing the hardware manufacturer into the role of trusted third party.
 In their model, the wallet would be a means of transacting with money held in custodial accounts that would be reconciled ex post.
 3.1 Fair exchange
-If we define a fair exchange as a transaction mechanism that ensures that transacting parties either each receive or each do not receive what it expects (and do not receive “any additional information about the other’s item”) [19], then it is not possible to have a fair exchange between two parties without the involvement of a third party; this is a foundational truth [20].
-With the exchange of physical objects, possession of the object itself is sufficient to demonstrate the veracity of the exchange that has taken place.
+If we define a fair exchange as a transaction mechanism that ensures that “at the end of an exchange [between two players], either each player receives the item it expects or neither player receives any additional information about the other’s item” [Asokan 1998], then it is not possible to have a fair exchange between two parties without the involvement of a third party; this is a foundational truth [Pagnia 1999].
+With the exchange of (unregistered) physical objects, possession of the object itself can be sufficient to demonstrate the veracity of the exchange that has taken place.
 But with digital objects the situation is more complicated.
-It is possible to have fair exchange without connection to the Internet, provided that there are third parties accessible locally, perhaps via a local network, as might be possible with ad hoc networks.
-Proponents of approaches requiring certified hardware have proposed embedding into end-user devices a tamper-resistant “secure element” that can be used to systematically restrict what the device can do, disclose the immutable identity of the device to a third party, or furnish information to third parties about what the device had observed or done.
-The secure element would be impenetrable to the user, and the user would not be able to modify its behaviour without rendering the device inoperable.
+It is possible to have fair exchange without connection to the Internet, provided that a mutually trusted third party is accessible locally, perhaps via a local network, as might be possible with ad hoc networks.
+In particular, it is not necessary for transacting parties to have connectivity to the issuer of the assets being transacted, since the mutually trusted third party might not be the issuer, and it is not necessary for transacting parties to have connectivity to providers of attributable accounts, such as banks, since the mutually trusted third party might not be an account provider or custodian.
+Transacting parties can provide assurance, for example, in the form of cryptographic proofs or signatures, that any applicable redemption and compliance obligations will be fulfilled.
+Proponents of approaches requiring certified hardware have proposed embedding into end-user devices a tamper-resistant “secure element” that can be used to systematically restrict what the device can do.
+Such restrictions can be used to provide the immutable identity of the device to a third party, or to furnish information contained within the device to third parties [IETF RFC 9334 2023, Bela 2024].
+The secure element would be impenetrable to the user, and the user would not be able to modify its behaviour without rendering the device unusable.
 In effect, such an element effectively acts as the third party by proxy, by embedding the will of the third party into the operation of a device carried by a user.
 3.2 The perils of trusted hardware
-In the case of digital currency, the secure element would presumably be used as an on-board authority that would prevent double-spending by maintaining a record of transactions made by the device and disallowing users from transacting tokens that have already been signed over to others.
-This design model in particular, and trusted hardware in general, introduces a variety of problems.
+In the case of digital currency solutions that rely upon a hardware root of trust, the secure element would presumably be used as an on-board authority that would prevent double-spending by maintaining a record of transactions made by the device and disallowing users from transacting tokens that have already been signed over to others.
+This design model in particular, and trusted hardware in general, introduces a variety of problems that can be avoided if support for dual-offline payments is not mandatory.
 We consider some of those problems here:
@@
 If the trusted hardware fails to function as advertised or is compromised by an attacker, then users could potentially transact the same tokens an arbitrary number of times without bound.
-The system-level costs of such an attack would be huge, and the incentives would be tremendously enticing for attackers.
-We note that modern devices, such as debit cards, that rely upon tamper-proof hardware for the exchange of value generally interact with accounts.
+Whereas with account-based systems, such as credit and debit cards, the cost associated with the compromise of a hardware device is limited to the size of the account with which it is associated, for a privacy-respecting system in which money is not linked to identity of the payer, the system-level costs of such an attack would be huge, and the incentives would be tremendously enticing for attackers.
+We note that modern devices, such as debit cards, that rely upon tamper-proof hardware for the exchange of value generally interact with accounts [ISO 8583 2023].
 If their security is compromised, then the value available to attackers is limited by rules and operational procedures implemented by custodians.
-Without custodial accounts, such backup mechanisms are not possible, and so, a successful attack on the hardware root of trust would imply that it would be necessary for all users to replace their devices.
-Thus, it is reasonable to conclude that it is preferable to rely upon a protocol to achieve the required level of security, as can be done when one of the parties has access to the network, rather than to expect that trusted hardware modules would guarantee that all transactions would take place within the security envelope of a trusted custodian.
-To assume that all hardware wallets must take the latter approach because some circumstances might require it is tantamount to overspecification at the expense of users who would readily hold assets offline but transact them offline.
-It is worth noting that when Chaum, Grothoff, and Möser argue that “any economically producible device that stores tokens with monetary value in an individual’s possession, and enables offline transactions… will be the target of successful forgery attacks…” [34], they are referring to wallets that perform dual-offline transactions, in contrast to wallets that hold tokens offline and transact them online.
+Without custodial accounts, such backup mechanisms are not possible, and so, the consequences of a successful attack on the hardware root of trust, such as one that impacts the cryptographic algorithm implemented in hardware or the mechanism that protects it from tampering, might not be addressable by a software upgrade, in turn requiring all users to replace their devices.
+Therefore, if both privacy and security are objectives of a digital currency system, it is preferable to rely upon a protocol, rather than the impermeability of end-user devices, to enforce the rules necessary to achieve the required level of security.
+Such a protocol can be used whenever one of the parties has access to the network, and under such circumstances, it would be better to allow the use of such a protocol rather than assume that trusted hardware modules would not be breached and would guarantee that all devices operate within the security envelope of a trusted custodian.
+To assume that all hardware wallets must take the latter approach because some circumstances might require it is tantamount to overspecification at the expense of users who would readily hold assets offline but transact them online.
+It is worth noting that when Chaum, Grothoff, and Moser argue that “any economically producible device that stores tokens with monetary value in an individual’s possession, and enables offline transactions… will be the target of successful forgery attacks…” [Chaum 2021], they are referring to wallets that perform dual-offline transactions, in contrast to wallets that hold tokens offline and transact them online.
+While overspecification of device requirements might certainly appeal to device manufacturers seeking compulsory mandates for their proprietary technology, it is important to consider that protocols that do not rely upon a hardware root of trust can function perfectly well outside the dual-offline context.
+Such protocols do not depend upon the imperviousness of end-user devices to attacks by those who possess them, and such attacks tend to fall within the capabilities of adversarial state actors and global crime syndicates.
   2. 2.
@@
 Because trusted hardware is impenetrable to users, users cannot know with certainty what their devices are doing.
-The inability to analyse the behaviour of a device undermines the trust that users can and should place in the device.
+The inability to analyse the operations of a device undermines the trust that users can and should place in the device.
 In essence, the device serves a second master that is not the user.
-Users cannot trust what they cannot verify, so they have every reason to be suspicious of unchecked promises on the part of hardware manufacturers or other privileged guarantors [21].
-For example, secure elements can furnish information that can be used to identify the user or the user’s previous transactions.
-By contrast, with open hardware, skilled members of the community can conduct audits and provide reports to the public.
+Users cannot trust what they cannot verify, so they have reason to be suspicious of unchecked promises on the part of hardware manufacturers or other privileged guarantors [Stallman 2002].
+For example, tamper-proof components might store and furnish information that can be used to identify the user or the user’s previous transactions.
+By contrast, with free hardware design111“design that permits users to use the design (i.e., fabricate hardware from it) and to copy and redistribute it, with or without changes” [Stallman 2015]., skilled members of the community, unbound by nondisclosure agreements, can conduct audits on the design of the hardware and provide reports to the public.
 Of course, users with the means to do so can personally conduct the audits themselves or hire contractors to conduct the audits for them.
-However, it is not necessary for every user to actually run an audit; evidence suggests that similar benefits are achieved so long as all users can perform audits if they choose to do so [22].
+However, it is not necessary for every user to actually run an audit; evidence suggests that similar benefits are achieved so long as all users can perform audits if they choose to do so [Heydt-Benjamin 2006].
   3. 3.
-    Certified hardware discourages innovation.
+    Certified hardware discourages beneficial innovation.
 Users who cannot create their own devices cannot innovate.
 Instituting a certification requirement for hardware serves to ensure that only those with the time and resources to pursue certification will be able to produce working devices.
-Regulations that increase the cost of entry are known to decrease entrepreneurship as well as innovation by existing firms [23].
-Ordinary users and hobbyists will not be able to develop their own devices, creating a barrier to public understanding the of the core technology, encumbering public debate and consideration of alternative designs [24, 25].
+Regulations that increase the cost of entry are known to decrease entrepreneurship as well as innovation by existing firms [Kozeniauskas 2022].
+Ordinary users and hobbyists will not be able to develop their own devices, creating a barrier to public understanding the of the core technology, encumbering public debate and consideration of alternative designs [Felten 2002, Samuelson 2016].
+For this reason, hardware certification requirements may implicitly promote security through obscurity, contrary to accepted practices in security engineering [Anderson 2001].
   4. 4.
-    Certified hardware distorts the market by privileging its manufacturers.
-The requirement to use trusted hardware effectively locks the hardware manufacturer in to a privileged position in the market, wherein prospective manufacturers require certification by some authority (either directly by the government or by a delegated authority such as a trade organisation), and concomitant lobbying and regulation can restrict free entry, increasing market concentration [26].
-Furthermore, even if all manufacturers could have certification at zero marginal cost, the fact that fabrication facilities carry high fixed costs means that the market is likely to be highly concentrated, with manufacturers capturing the surplus [27].
+    Certified hardware undermines competition by privileging its manufacturers.
+The requirement to use trusted hardware effectively locks the hardware manufacturer in to a privileged position in the market, wherein prospective manufacturers require certification by some authority (either directly by the government or by a delegated authority such as a trade organisation), and concomitant lobbying and regulation can restrict free entry, increasing market concentration [Gutierrez and Philippon 2019].
+In essence, enforcing hardware certification transforms a market for consumer devices into something more closely resembling public infrastructure procurement.
+Furthermore, even if certification were not required or if all manufacturers could have certification at zero marginal cost, the fact that fabrication facilities carry high fixed costs means that the market is likely to be highly concentrated, with a few dominant manufacturers capturing the surplus [Gao and Kehrig 2017].
+Without a market to serve their needs directly, end-users will have no choice but to accept the solution that the government chooses to trust.
   5. 5.
@@
 But unlike trusted third parties on the network, users bind themselves to long-term relationships with personal devices that contain trusted hardware; the relationships are not much different between account-holders and custodians, wherein account-holders are subject to satisfying the requirements of custodians.
-The cost of switching bank accounts is generally only a matter of time and effort on the part of account-holders, whereas the cost of switching devices generally includes per-device manufacturing and retailing cost plus taxes and costs associated with physical acquisition and disposal of the devices.
+Although in principle, this can be regulated, implementation inexorably depends upon how custodians implement regulatory compliance and is subject to the limits of the effectiveness of regulation.
+The cost of switching bank accounts is generally only a matter of time and effort on the part of account-holders and nevertheless remains a notoriously difficult problem for regulators, whereas the cost of switching devices, such as smartphones, generally includes per-device manufacturing and retailing cost plus taxes, as well as costs associated with physical acquisition and disposal of the devices.
 In addition, while it is possible to implement both technical protocols and regulations that rely upon locally-trusted third parties, economies of scale mean that hardware manufacturers are likely to have global reach.
 It is reasonable to conclude that relationships between users and their devices featuring specific trusted hardware will be sticky indeed.
-For these reasons, we are unconvinced that trusted hardware solutions are appropriate as requirements for the design of tools that allow users to possess and control their own assets.
+For these reasons, we are unconvinced that trusted hardware solutions are appropriate as requirements for the design of tools that allow users to possess and control their own digital assets.
 At the same time, we would not suggest that trusted hardware should be forbidden.
-In some fraction of cases in which physical cash is not an option and network connections are not available, trusted hardware might be an alternative kind of third party for transacting parties to consider, although this possibility is not sufficient justification to institute a requirement.
-In all cases, both sender and recipient must willingly choose to trust the third party to consummate the transaction, and given the final three points in the last above, it is difficult to imagine that a user preference for a particular device manufacturer would be real.
-Consider, for example, that existing users of secure hardware in the form of electronic payment cards generally have little say in which payment network their issuing banks choose.
+In some fraction of cases in which physical cash is not an option and network connections are not available, trusted hardware might be an alternative kind of third party instrument for transacting parties to consider, although this possibility is not sufficient justification to institute trusted hardware as a universal requirement.
+In all cases, both sender and recipient must willingly choose to trust the third party to consummate the transaction, and given the final three points in the list above, it is difficult to imagine that users would be able to exercise free choice of whom to trust rather than be forced to accept a particular hardware manufacturer.
+Consider, for example, that existing users of secure hardware in the form of electronic payment cards often have little say in which four-point payment network their issuing banks choose, and even when they do, the set of existing card payment networks remains well-entrenched.
+Similarly, secure hardware solutions have implications for interoperability: If the payer and payee must use the same third-party hardware manufacturer, then network effects might tend to entrench a single platform or small set of platforms.
 3.3 Standards and lobbying
@@
 But this argument is unsound: Cash is an excellent offline payment method, and to our knowledge, few if any central banks have suggested that the purpose of CBDC is to replace cash outright.
-The argument implicitly equivocates between a concept of “offline” that denotes transactions wherein a user is not required to carry a device with network connectivity and a concept of “offline” that involves two counterparties physically meeting in a location without third parties or network infrastructure of any kind nearby.
-The former scenario is the only one in which users commonly use certified hardware today.
-The idea that cash could suddenly become unavailable for use in the latter scenario is an overly simplistic justification for hardware manufacturers to tout their products.
+The argument implicitly equivocates between a concept of “offline” that denotes a transaction wherein a user is not required to carry a device with network connectivity, versus a concept of “offline” that denotes a transaction between two counterparties that physically meet in a location without third parties or functioning network infrastructure of any kind nearby.
+The former scenario is the only one in which users commonly use certified hardware today, for example, single-offline POS payments (e.g.
+Apple Pay) and network applications that rely upon DRM (e.g.
+video streaming).
+Even the specialised Suica card offered by Japan Rail for use at its payment gates, which uses embedded trusted hardware to facilitate payments when connectivity to a central server is broken [Shiibashi 2008], could potentially use locally reachable points of trust to aggregate transaction information rather than assume that network connectivity is either all or nothing.
+Moreover, the idea that cash could suddenly become unavailable for general use in offline situations (the latter scenario) is an overly simplistic justification for a system-level requirement.
 We note that secure elements can be used for a variety of purposes, including the prevention of exfiltration of private data from a device.
@@
 “remote attestations”) that a device is operating within the security envelope of a trusted third party, whether a hardware manufacturer, service provider, or authority, implies a specific model for securing transactions that is not necessary for single-offline payments.
-For the reasons described above, we argue that single-offline payments would be both more robust and more secure if counterparties can trust network service providers, and the various methods network service providers can use to demonstrate the veracity of their claims, rather than having no choice but to rely upon trusted hardware physically held by the other counterparty.
-And so, for single-offline payments, trusted hardware adds nothing, but imposes significant costs, and therefore it must not be considered a requirement.
-Dual-offline transactions can certainly make use of remote attestation if both parties agree, but such transactions should not be seen as a general case for digital currency transactions, particularly given pervasive connectivity and the security benefits of using protocol-based security rather than introducing a requirement for counterparties to trust each other’s devices.
-For this reason, the development of standards that recommend or require hardware wallets to use a secure element for remote attestation are certainly problematic.
-Standards that recommend hardware wallets to use a secure element to prevent the exfiltration of private data might be appropriate in the case of transactions involving smartphones, because of their complexity and large attack surface.
-However, it is inappropriate to assume that all or most digital wallets will be smartphones, not least because of their intrinsic security and privacy risks, despite being the “benchmark in user-accessible features”, as Barresi and Zatti argue [5].
+For the reasons described above, we argue that single-offline payments would be both more robust and more secure if their locus of trust can be a network service, relying upon the various methods network service providers can use to demonstrate the veracity of their claims, rather than having no choice but to rely upon the integrity of trusted hardware physically held by the other counterparty.
+locus of trust network service certified hardware
+supports secure single-offline payments yes yes
+supports secure dual-offline payments no yes
+requires wallets to be linked to attributed accounts no yes
+allows anonymous payers (bearer instruments) yes no
+allows unregistered or uncertified end-user devices yes no
+devices may be built by user communities without approval yes no
+devices may be fully auditable by end-users yes no
+Table 2: Features of a network-based trust locus versus a hardware-based trust locus.
+Table 2 summarises the key differences between using a network service as a root of trust and using certified hardware as root of trust.
+Although certified hardware has the advantage of enabling dual-offline payments, such payments are not actually common in the digital economy.
+Secure network services can provide the same degree of transaction assurance that secure hardware devices can provide, with the only limitation being that they must be reachable by one of the transacting parties.
+At the same time, we have argued that because it is feasible for adversarial state actors or organised crime syndicates to compromise individual devices in their possession, it would be too much of a risk for a system relying upon such devices to allow them to spend money that is not linked to attributed accounts, because compromised devices would be able to reuse anonymous money (bearer instruments) without limit.
+However, with a network service, there is no system-level risk if an end-user device is compromised, since the locus of trust is elsewhere.
+A network service that provides payment finality222the state in which obligations are discharged between transacting parties.
+for bearer instruments can be operated by a trusted third party, as in GNU Taler, or it can be oblivious and based on public commitments, so that compromising it does not enable double-spending [Friolo 2025].
+And so, for single-offline payments, trusted hardware adds nothing, but imposes significant costs and limitations, and therefore it must not be considered a system-level requirement.
+Dual-offline transactions can certainly make use of remote attestation if both parties agree, but such transactions should not be seen as a general case for digital currency transactions, particularly given pervasive connectivity and the security benefits of using protocol-based security with an external root of trust rather than introducing a requirement for counterparties to trust each other’s devices.
+For this reason, the development of standards that recommend or require hardware wallets to use a secure element for remote attestation are certainly problematic, while perfectly functional wallet devices that provide better end-user affordances could be designed to strictly facilitate single-offline payments instead.
+Such devices do not necessarily require network capabilities themselves, provided that their transaction counterparties are online.
+Standards that recommend hardware wallets to use a secure element to prevent the exfiltration of private data might be appropriate in the case of transactions involving smartphones, because of their complexity and exposure to malware attacks.
+However, it is not necessarily the case that all or most digital wallets will be smartphones, despite being the “benchmark in user-accessible features”, as Barresi and Zatti have argued [Barresi 2023].
+One reason is price: Although virtually everyone must have a way to engage with the economy, a substantial share of the population cannot afford a smartphone.
+In 2024, the global median price of a smartphone was US\$92.59, representing 11% of global median per capita income [GDIP 2024].
+Another reason is security: with ever more complex operating systems and applications, smartphones are increasingly plagued by security vulnerabilities [NIST], many of which specifically target banking credentials [Fadilpašić 2025].
+For a sensitive function like payments, it might be safer and more appropriate to use a dedicated device [Bizama 2025].
 4 Digital currency means tokens, not accounts
-Barresi and Zatti assume that wallets would represent end-user money as account balances [5].
-However, for ordinary consumers and non-financial businesses to hold digital currency directly, there must be some way for digital currency to be held outside accounts.
-Accounts generally involve a relationship between a beneficial owner (the account holder and a fiduciary, typically a regulated financial institution.
-A salient property of accounts is that access to assets held within accounts is determined by the identity of the account holder [28].
+Barresi and Zatti assume that wallets would represent end-user money as account balances [Barresi 2023].
+However, for ordinary consumers and non-financial businesses to hold digital currency directly, there must be some way for end-users to hold digital currency without the requirement to have an attributed account.
+Accounts generally involve a relationship between an account holder and a fiduciary, typically a regulated financial institution.
+A salient property of accounts is that access to assets held within accounts is determined by the identity of the account holder [Auer and Böhme 2020], and the balance of an account represents an entitlement that can be adjudicated, for example, in a court of law.
 In contrast, tokens are assets that can be held outside of accounts.
-Tokens can be bearer instruments, meaning that they can be used in a way that does not index the identity of their owners.
+Tokens can be bearer instruments, meaning that they can be used in a way that does not reference the identity of their holders.
 Payments can be held offline and transacted online, so the user experience would be similar to an EMV card payment except in which the payer presents digital assets rather than instructions to debit an account.
-Compared to assets held within accounts, assets that are held directly by their owners have a different set of security requirements for their owners.
-For example, owners of assets held within accounts can rely upon various identification mechanisms, such as PIN numbers and two-factor authentication, to protect their assets if their cards or mobile devices are stolen, and asset custodians can also provide credit to account holders, such that if assets are spent unlawfully, the account holders can ask the asset custodians cover the value of their losses.
+Compared to assets held within accounts, assets that are held directly by their holders have a different set of security requirements.
+For example, holders of assets held within accounts can rely upon various identification mechanisms, such as PIN numbers and two-factor authentication, to protect their assets if their cards or mobile devices are stolen, and asset custodians can also provide assurances to account holders that if assets are spent against their wishes, then the account holders can ask their asset custodians cover the value of their losses.
 In contrast, the bearer is responsible for bearer instruments.
-If cash is lost or stolen, owners generally have little recourse beyond law enforcement mechanisms to recover their money, and such mechanisms often fail to recover the value.
+Persons whose cash is lost or stolen generally have little recourse beyond law enforcement mechanisms to recover their money, and such mechanisms often fail to recover the value.
 The potential for loss or theft exists not only physically but also electronically.
@@
 The prevailing model for handling devices that no longer work or have been compromised is to wipe their contents completely or to replace them wholesale.
-When such devices are used to access accounts, such replacement is not much of a problem, although bearer instruments inexorably expose their owners to new risks that require different solutions.
-Some central banks, including the ECB [14], the US Federal Reserve [29], and the Bank of England [31], have presented arguments that holding limits for wallets are necessary for financial stability and to prevent bank runs.
-Such arguments are nonsense.
-Limiting CBDC withdrawals, rather than aggregate holdings, would be no less effective in preventing bank runs, and the approach of limiting bank withdrawals has been used successfully in the past, notably in Greece in 2015 [30].
+When such devices are associated with accounts, replacing devices that have been demonstrated to be insecure might be operationally expensive or cumbersome, but the actions of a compromised device are ultimately traceable to the account with which it is associated, thus giving the account holder ways to mitigate the effects.
+Although the potential for loss or theft of bearer instruments inexorably constitutes a risk that requires a different kind of solution, approaches well-suited to digital technology, such as backup copies and withdrawal records that can be used by victims of theft to selectively de-anonymise their own tokens that had been stolen, can be robust and effective in mitigating such risks.
+Some central banks, including the ECB [European Commission 2023], the US Federal Reserve [Federal Reserve 2022], and the Bank of England [Bank of England 2023], have presented arguments that holding limits for wallets are necessary for financial stability and to prevent bank runs.
+However, as an alternative to the heavy-handed approach of monitoring or restricting wallets, an approach of limiting CBDC withdrawals from banks would be no less effective in preventing bank runs or bank disintermediation than limiting cash withdrawals from banks, and the approach of limiting bank withdrawals from banks has been used successfully in the past, notably in Greece in 2015 [Associated Press 2018].
+In this sense, CBDC is no different from cash.
+To argue that requiring wallets to be managed by parties other than their holders is necessary to address financial stability problems resulting from reduced liquidity is to conflate the technical design requirements for a wallet with economic justifications for eliminating the ability for individuals to hold money outside accounts.
+However, although some banks and governments might see value in eliminating cash-like instruments and might view CBDC as a vehicle for achieving that objective, such an argument is not commonly articulated.
 Importantly, limiting CBDC withdrawals can be done without requiring wallets to implement account-like control mechanisms that undermine the will of their users.
-Barresi and Zatti suggest that wallets could perform “behavioural analytics” on their users [5]; little could be more disrespectful of a user’s autonomy.
-Although a third party is necessary for fair exchange (and indeed trusted hardware or closed-source software blobs can be this third party de facto), Pagnia and Gärtner have shown that the role of the third party can be limited, even zero-knowledge: tokens can be transferred without the third party knowing what is being transferred or having a custodial relationship with either party [20].
+Barresi and Zatti suggest that wallets could perform “behavioural analytics” on their users [Barresi 2023]; little could be more disrespectful of a user’s autonomy.
+Similarly, the argument that digital currency could be used as a way to implement “programmable money” that can only be spent in pre-defined ways is a non-starter.
+Such a mechanism would not only undermine the singleness of money but also open a Pandora’s box of ethical concerns [Li 2025], some of which have already been realised by the failed cashless debit card scheme in Australia [Greenacre 2023].
+However, banks could impose simple rules to limit the aggregate size of withdrawals that can be made from their accounts into CBDC, both in a single transaction and over a period of time, and such rules could be potentially combined with identity information to limit the amount that an end-user could withdraw throughout the system.
+Such rules could also be combined with periodic token expiry, as a mechanism to prevent hoarding [Goodell 2021].
+Periodic token expiry is less intrusive than mechanisms that restrict or control wallets directly, and it can be made transparent to users by implementing a mechanism for wallets to automatically refresh any assets that are close to the expiration time.
+Although a third party is necessary for fair exchange (and indeed trusted hardware or closed-source software blobs can act as this third party de facto), Pagnia and Gärtner have demonstrated that the role of the third party can be performed “in zero-knowledge”.
+Specifically, a third party can facilitate final settlement of an asset transfer involving two transacting parties, without having a custodial relationship with either party, without knowing the identity of either party, without knowing how much is being transferred, and even without knowing that tokens are being transferred at all [Pagnia 1999].
+Importantly, with oblivious management, an issuer does not rely upon a central ledger to track tokens individually, although the issuer must trust that the system that collects transactions and produces rolled-up aggregation entries, one per time period, will not equivocate.
+This trust can be enhanced by requiring that system to externalise its commitment to a single history, for example by sending some or all of its aggregation entries to the issuer, or by writing them to a public bulletin board or distributed ledger.
 The idea that digitalisation implies the use of accounts is not true in general.
-Digital tokens can be real indeed, with ownership demonstrable by transferring it: using a single-use private key to transfer the asset to another single-use private key.
+Digital tokens can indeed be held directly as data, with ownership demonstrable by transferring it: using a single-use private key to transfer the asset to another single-use private key.
+A transaction is a cryptographically signed mapping of that single-use key to a commitment to assign the asset to another single-use key, and the “proof of provenance” offered by an integrity provider demonstrates that the first single-use key has been irreversibly mapped in this way.
+Users will require a mechanism for generating keys, but open source tools plus a good source of entropy are sufficient for this.
 The requirement of an account in the form of a “wallet provider” is not true for cryptocurrencies in general, nor does a private key constitute a proof of identity when it is created to be used exactly once.
 It is true that many Bitcoin users rely on service providers to manage their holdings.
-However, to argue from this that self-custody of tokens will never take off is pure speculation, and it ignores the fact that most Bitcoin users are speculators with no expectation that they would use Bitcoin for ordinary retail purposes.
-Many people carry around an amount of cash that is useful for retail purchases but small enough that they could lose it without too much regret, despite the fact that losing it is forever.
-Digital assets can certainly be held by users without accounts of any kind, for example, as a payer can in Chaum’s model.
+However, to argue from this that self-custody of tokens will never take off is pure speculation.
+Many people deliberately carry around an amount of cash that is useful for retail purchases but small enough that they could lose it without too much regret, despite the fact that losing it is forever.
+They exercise a free choice to determine how much cash to carry, without being forced via technology to adhere to some limit.
+More importantly, even those who might be comfortable with a limit might be uncomfortable with the mechanism used to impose the limit, particularly if it requires them to use a device controlled by someone else (as described in Section 3.2).
+Digital assets can indeed be held by users without accounts of any kind, just like cash and other physical assets.
+An example of a digital asset that can be held without an account is an e-cash token in Chaum’s model [Chaum 1982, Chaum 2021], which requires every transaction to function as a redemption, and it has also been shown that digital assets can be transferred without redemption [Goodell 2023, Friolo 2025].
 Accounts for recipients might be desirable, but they are not technically necessary.
-The involvement of a third-party to prevent double-spending does not imply that the third party must have a role as custodian or account provider, or knowledge of the asset being transferred, the circumstances of the transfer, or the identity of either party.
+As described above, it has been proven that the prevention of double-spending does not imply the existence of a traceable record of all the holdings and movements through sequential time.
+In Chaum’s original proposal, an asset can be transferred to a recipient without revealing information about the payer, and, unlike oblivious management, double-spending is prevented by maintaining a record of the spent tokens.
+In the general case, and in the approach taken by oblivious management, to prevent double-spending, it is sufficient to register non-repudiable commitments from payers, which may take the form of a hash or zero-knowledge proof that can be verified without knowledge of the transaction [Pagnia 1999], and proofs of these commitments can be delivered along with the asset data rather than centrally recorded [Goodell 2023, Friolo 2025].
+Oblivious management presupposes that issuers can be held responsible for what they have issued and that proofs of commitments that cryptographically demonstrate the validity and integrity of an asset constitute sufficient evidence of an issuer’s responsibility to honour a redemption.
+As discussed earlier, specific payment architectures that implement these mechanisms have been proposed.
 Tokens are also important for privacy.
-Systems that make use of transactions involving balances, even if the transactions themselves are privacy-preserving [36, 37], intrinsically require end-users to retain data produced during a transaction that can be used to prove that the transaction was carried out according to the rules.
+Systems that make use of transactions involving balances, even if the transactions themselves are privacy-preserving [Wüst 2022, Kiayias 2022], intrinsically require end-users to retain data produced during a transaction that can be used to prove that the transaction was carried out according to the rules.
 This means that transactions are not independent of one another.
-The property of transaction independence is important because retaining data related to past transactions introduces the risk that the user could be de-anonymised via poor operational security or blackmailed to reveal information about those transactions [38].
-Finally, the idea that digital currency intrinsically requires accounts is not true; central banks can certainly issue tokens that exist outside of accounts, just as they do with cash, and a token like this would be a liability on a central bank’s balance sheet, not unlike cash or central bank reserves.
+The property of transaction independence is important, because retaining data related to past transactions introduces the risk that the user could be de-anonymised via poor operational security or blackmailed to reveal information about those transactions [Friolo 2025].
+Because it must be possible for balances to be tracked through all of the movements of that balance, transactions in systems that represent money as balances cannot achieve transaction independence.
+However, systems that implement money as tokens are not subject to this constraint.
+Finally, the idea that digital currency intrinsically requires accounts is not true; central banks can certainly issue tokens that exist outside of accounts, just as they do with cash, and a token like this would be a liability on a central bank’s balance sheet, not unlike cash.
 5 Separating issuance from transaction processing
-Even if transactions are done with network access to a mutually trusted third party, there is no reason to assume that the issuer must be involved in transaction processing in real-time.
-Chaum’s original proposal assumed such a dependency [33], but this is not a requirement in general.
-In particular, it is possible to register a (zero-knowledge) commitment to transfer an asset without communicating with the issuer, and for a transaction to have finality it is sufficient for both parties have faith in a mutually trusted third party to eventually post the commitment to a bulletin board recognised by the issuer [10].
-Also, if tokens carry their own provenance information, then there is no need for the issuer to track tokens that it had seen, either [10].
-It is certainly possible to transact money issued by a central bank in a decentralised system, and wallets can be expected to respect this possibility.
-Trusting a central bank (or other issuing authority, such as those that issue e-money) to issue money and maintain its integrity is certainly different from trusting it to process transactions.
+Even if transactions are done with network access to a mutually trusted third party, there is no reason to assume (a) that the third party must be the issuer, or (b) that the issuer must be involved in transaction processing in real-time.
+Chaum’s original proposal assumed both conditions [Chaum 1982], but they are not requirements in general.
+In particular, with oblivious management, it is possible to register a (zero-knowledge) commitment to transfer a digital asset without communicating directly with the issuer.
+For a transaction to have finality, assuming every transfer specifies the payee with a unique single-use key, it is sufficient for both parties have confidence in a mutually trusted third party, acting as a notary, to eventually post the commitment (or an aggregation that comprehends the commitment) to an append-only bulletin board that maps single-use keys to commitments and that is trusted by the issuer [Goodell 2023].
+The append-only bulletin board can be operated by the issuer directly, by a party that the issuer trusts, or as a distributed ledger operated by a set of parties that the issuer trusts in aggregate.
+Provided that the operator of the bulletin board does not equivocate, then since no two tokens can be transferred to the same single-use key and have a valid proof of provenance, it is not necessary for the issuer to maintain a ledger that records the registration status of individual tokens [Goodell 2023].
+Even if we assume that all tokens are issued by a central issuer, such as a central bank, it is certainly possible to use a decentralised system to conduct transactions using those tokens, and wallets can be designed to respect this possibility.
+Trusting a central bank (or other issuing authority, such as one that issues e-money) to issue currency and maintain its integrity is certainly different from trusting it to process e-cash transactions.
 In particular, it is possible to use a decentralised system to transact centrally issued money.
-The use of a secure element to establish trust implicitly requires the transacting parties to trust a hardware manufacturer or other central authority to facilitate transactions.
-Such a role for a unified central authority might be incompatible with the concept of a decentralised system.
-Finally, requirements for hardware wallets used for digital currency should not be needlessly inconsistent with requirements for hardware wallets used for other digital assets.
-It is also worth considering that all users cannot be expected to trust the same trust anchors equally.
-Although the fact that assets held in offline non-custodial wallets can be transacted online is sufficient to reject the requirement of a hardware root of trust for security, even the case of dual-offline transactions poses a problem for requiring the use of secure elements.
-Specifically, there is the matter of who the transacting parties choose to trust.
-We imagine that parties choosing to transact online could potentially choose from a variety of manufacturers of hardware roots of trust based upon their needs and security models, whereas forcing manufacturers of hardware wallets to embed hardware certified by a specific authority, with the expectation that this authority is universally and equally trusted by all users, is not appropriate.
-We believe that allowing users to retain a choice of mutually trusted hardware security solutions, such as removable eSIM devices, is appropriate, although it might be impractical for users to carry around as many hardware devices as there are commonly trusted third parties with their various counterparties.
-The fact that hardware devices impose a trust model based upon their specific manufacturer is yet another reason to reject the idea that such devices could function as a public utility.
+Just as cash can be managed with a two-tiered approach, e-cash can also be managed with a two-tiered approach.
+If we assume, following Chaum, that individual e-cash assets are generated for one-time use by consumers, then a central bank can provide private-sector banks with “genesis” assets that can be burned, pari passu, by private-sector banks in the process of creating single-use e-cash tokens for consumers.
+Then, the single-use assets can be recirculated within the system by private-sector banks indefinitely via the same process before ultimately being redeemed with the central bank [Friolo 2025].
+The use of a secure element to assure trust seems to implicitly require the transacting parties to respect the authority of a specific hardware manufacturer, and the power implicit to this authority might be incompatible with the concept of decentralisation if all transacting parties are forced to use secure elements produced by a limited set of manufacturers.
+To require the use of a hardware root of trust with remote attestation unnecessarily limits decentralisation.
+The fact that assets held in offline non-custodial wallets can be transacted online is sufficient to reject the requirement of a hardware root of trust for security, because trusted hardware is unnecessary for online and single-offline payments (see Table 2), and not every transaction must be treated as a dual-offline transaction.
+However, even if we accept that trusted hardware is necessary for dual-offline payments, a key problem remains: Which hardware manufacturer should transacting parties choose to trust?
+The risk implicit to requiring all users of a system to respect the same set of trusted authorities is well-documented [Kushner 2013].
+The incentives for operators of infrastructure, such as telecommunications providers operating under public monopoly, to discourage beneficial innovation are also well-documented [Tran 2019].
+We imagine that parties choosing to transact online could potentially choose from a variety of trust providers based upon their needs and security models, whereas forcing manufacturers of hardware wallets to embed hardware certified by a specific authority, with the expectation that this authority is universally and equally trusted by all users, may be too demanding.
+We believe that allowing users who need dual-offline payments to retain a choice of mutually trusted hardware security solutions, such as removable eSIM devices, is appropriate, although it might be impractical for users to carry around as many hardware devices as there are commonly trusted third parties with their various counterparties.
+Interoperability does not solve this problem, since it is the hardware itself that must be trusted.
+Substituting consumer choice with public infrastructure might seem like a tempting solution, although it entails requiring consumers to relinquish some of their autonomy.
+Finally, there is no reason for which the technical requirements for hardware wallets used for digital currency in general should be inconsistent with the technical requirements for hardware wallets used for other transferable digital assets, including those that do not require a hardware root of trust because they are not expected to support dual-offline payments.
+Both kinds of wallets, those that satisfy special requirements to support dual-offline payments and those that do not, should be able to work within a digital currency system.
 6 Wallets as containers for bearer instruments
+possession self-custody versus third party custody
+ID requirement yes (account) versus no (bearer)
+proximity in-person (POS) versus remote (e-commerce)
+device authorisation certified hardware versus public protocol
+form physical (cash) versus digital
+issuer central bank versus private sector entity
+regulation bank money versus e-money
+Table 3: Some dimensions by which payment technologies can be described.
 For a set of specific technical requirements for digital currency in general to be credible, it must address the requirements of its various stakeholders.
 One of the unfortunate characteristics of the current debate is the diversity of conflicting assumptions about the nature of payments and the relative importance of different kinds of payment scenarios.
-In reality, there are many different ways that people make payments today: with and without custodians, locally and at a distance, with and without electronic devices, with and without money issued by a central bank.
-Money is transferred to pay invoices, to purchase goods and services, to give gifts, to pay taxes, or to reallocate funds within an organisation or family.
+In reality, there are many different dimensions that can be used to describe how people make payments today: with and without custodians, locally and at a distance, with and without electronic devices, with and without money issued by a central bank; see Table 3.
+Money is transferred with finality333as a way of discharging obligations between transacting parties.
+for a variety of reasons, including to pay invoices, to purchase goods and services, to give donations, to pay taxes, to repay loans, or to transfer funds within an organisation or family.
 These different scenarios imply a plethora of different costs and risk profiles: Buying a cup of coffee is different from buying a house.
 We should accept that there will not be a single payment system or key mechanism to rule them all, although it might be possible to develop a common framework that can accommodate a variety of different payment systems.
-The case for trusted hardware, as for other core mechanisms of proposed payment systems, involves the requirement for assessing the integrity of money in a fair exchange scenario wherein counterparties lack access to a mutually trusted third party.
-In principle, trusted hardware can act as a trusted third party to enforce system-level rules on user devices directly, by implementing rules that prevent assets from being spent more than once.
+The case for trusted hardware, as for other core mechanisms of proposed payment systems, involves the requirement to ensure the integrity of money when counterparties lack access to a mutually trusted third party.
+In principle, trusted hardware can act as a trusted third party to enforce system-level rules on user devices directly, by implementing rules that prevent digital assets held in a wallet from being spent more than once.
 However, considering the security risks associated with compromised devices, it is better to enforce such rules over the network instead, if it is possible to do so.
-This can be done, for example, by invoking a trusted third party to validate all transactions, as proposed by Chaum and others [33, 34].
-However, there is reason to believe the costs, risks, and public accountability concerns associated with having a central operator dictate that history unchecked are too great for the CBDC use case [35], and the most important costs and risks would be borne not only by the users but also by the central operator.
-A modern alternative involves using a distributed ledger system to put the money itself on a blockchain, the way that UTXO designs (e.g.
-Bitcoin [32], Monero [39]) and state-transition designs (e.g.
-Ethereum [40]) do.
-Such approaches are unwieldy because they require the ledger to track all of the tokens individually, and transaction finality requires global consensus.
+This can be done, for example, by requiring the issuer to validate all transactions, as proposed by Chaum and others [Chaum 1982, Chaum 2021].
+However, there is reason to believe the costs, risks, and public accountability concerns associated with having a single central operator perform issuance, settlement, and compliance functions might be too great for the CBDC use case [Goodell 2021].
+A modern alternative to the centralised model involves using a distributed ledger system to put the money itself on a blockchain, the way that UTXO designs (e.g.
+Bitcoin, Monero [Van Saberhagen 2013]) and state-transition designs (e.g.
+Ethereum [Buterin 2013]) do.
+However, such approaches are unwieldy because they require the ledger to track all of the tokens individually, and transaction finality requires global consensus, introducing a system-level economic externality with every transaction, exposing users to significant management costs, and, in some cases, creating uncertainty about finality.
 In practice, most cryptocurrency users rely upon custodial accounts, negating a foundational benefit of digital currency.
-However, distributed ledgers and blockchains can be used to support and enable payment systems in a variety of ways, including by providing non-repudiable evidence that a transaction had taken place, even if an individual service provider might equivocate [41].
-The digital currency design proposed by Goodell et al [10] and further elaborated by Friolo et al [38] addresses the scalability problems associated with modern cryptocurrency systems while strongly preventing double spending without trusted hardware.
-In particular, it avoids requiring the issuer to participate in every transaction of its tokens, it avoids requiring system operators to manage a list of tokens that have been created or spent, and it avoids requiring users to have access to a global network at the time of the transaction.
-Any third party chosen by the spender and accessible to the recipient at the time of the transaction can suffice.
-Proposed solutions that rely upon stateful assets and oblivious transfers can work technically without distributed ledgers, but in the end, digital currency systems require a means to support an immutable, shared history mutually accepted among independent actors, which is the salient characteristic of a distributed ledger [42].
-The immutability of a distributed ledger is emergent property of the independence of the participants.
-So, although using a distributed ledger to record and process tokens directly on the ledger might bring unacceptable costs in terms of scale, efficiency, and operational burden, we should not rule out using distributed ledgers for what they do best: externalising commitments so that parties responsible for maintaining the history do not equivocate [10].
-These commitments can be aggregated in such a way that the ledger grows at a fixed rate determined in advance, and each transaction carries zero marginal cost to the ledger operators.
+feature UTXO (e.g. Bitcoin) Chaum (GNU Taler) Oblivious management
+assets represented directly on a ledger when issued yes no no
+assets represented directly on a ledger when spent yes yes no
+bundled functions shared across network peers yes no no
+finality assurance supported by network peers yes no optional
+issuance separate from settlement and compliance no no yes
+settlement separate from compliance no no yes
+Table 4: Features of selected token-based payment systems.
+And so, instead of aiming to share the responsibility for a bundled set of issuance, settlement, and compliance enforcement functions across a set of peers, it might be better to decompose the functions into distinct roles that can be performed by different parties, enabling the costs and risks to be allocated more efficiently.
+Decentralised infrastructure can be used to support and enable payment systems in a variety of ways, including by providing non-repudiable evidence that a transaction had taken place, even if an individual service provider might equivocate [Goodell 2026].
+Table 6 provides a comparative overview of the various token-based payment systems that we have mentioned in terms of their key features.
+The digital currency design proposed by Goodell et al [Goodell 2023] and further elaborated by Friolo et al [Friolo 2025] offers an alternative approach to achieving scalability, replacing global consensus about the status and ownership of individual assets, and the related costs of such consensus common to modern cryptocurrency systems, with a combination of local trust and a mechanism for keeping local trust providers honest.
+By integrating a fungible token created using Chaum’s method that delivers privacy by design for consumers into a non-fungible asset that can be transferred with strong traceability following the initial creation, it enables the separation of the costs and risks associated with the aforementioned roles that enable payments.
+With this design, the third parties that provide transaction finality do not manage tokens, the issuer is not involved in individual transactions, and regulated entities, such as banks, can continue to provide compliance enforcement (e.g.
+income monitoring for the purposes of taxation and sanctions enforcement) as a condition of exchanging a spent token for a fresh token or another form of money.
+Oblivious management can work technically with a centralised notary, but in the end, digital currency systems require common agreement about history, which is the salient characteristic of a distributed ledger [ISO 22739 2024].
+Agreement about history does not imply shared knowledge of the holdings or transfers of individual assets.
+If assets maintain their own provenance information with references to ledger state, then having an integrity provider record only the zero-knowledge commitments of payers to their transactions, and having the ledger record only the rolled-up aggregation of such commitments, is sufficient.
+These commitments can be aggregated in such a way that the ledger grows at a fixed rate determined in advance, without individual transactions introducing system-level externalities, and such that each transaction carries zero marginal cost to the ledger operators.
+Some established DLT systems, such as Arbitrum, already aggregate transactions in this way [Kalodner 2018].
+Proof that a transaction is part of a fixed-size aggregation, which, subject to foundational cryptographic assumptions about the uniqueness of hashes, may be comprehensive of an arbitrary number of transactions, is sufficient.
 Trusted computing is not necessary to protect users from loss or theft, either.
@@
 Users can also upload copies of their tokens to non-custodial service providers, such as cloud storage services.
+As described in Section 4, double-spending is prevented by the requirement that every asset is bound to a specific integrity provider and a specific single-use key for specifying the next transaction on the asset.
 A much better way to address the limited set of cases in which “fully offline” payments (that is, those where neither the payer nor the payee have access to any suitable third parties or network infrastructure) are necessary would be to support the use of cash itself where cash continues to work well, such as in remote locations and environments with low connectivity, while avoiding all of the aforementioned problems associated with trusted hardware.
-To address the reduction of cash in the “built” environments by providing digital currency, whilst encouraging the provision of facilities for converting between cash and diital currency at the boundary of “unbuilt” environments, would be an entirely legitimate approach.
-In particular, if one potential benefit of CBDC in an age of deteriorating cash infrastructure is the ability to electronically transport cash to remote population centres during crises [43], then there is no reason to require that central bank money distributed in this manner must remain digital.
-For example, it can be exchanged for physical tokens specific to that population centre, much as cash and bank money are exchanged for chips at a casino.
+To address the observed reduction of cash acceptance in urban environments [Mansuri and Vyas 2024, De Nederlandsche Bank 2025] by providing digital currency, whilst encouraging the provision of facilities for converting between cash and digital currency at the boundary of “unbuilt” environments, would be an entirely legitimate approach.
+In particular, if one potential benefit of CBDC in an age of waning cash infrastructure [^4] Evidence for the decline of cash is growing [Access to Cash Review 2019, Tisher 2020], and it is acknowledged as an important risk by central banks, including the ECB [European Commission 2023a].
+is the ability to electronically transport money to remote population centres during crises [Goodell 2020], then there is no reason to require that central bank money distributed in this manner must remain digital.
+For example, service providers in a particular locality can provide physical tokens in exchange for CBDC, where the CBDC is transferred using one of the methods described earlier, much as a casino can provide chips in exchange for cash or bank money.
 None of this is to say that we should forbid limited groups of users from trusting certified devices, if that is what they freely choose, to conduct transactions without network connectivity to a third party.
@@
 However, at a system level, requiring all users to purchase and carry certified devices, even if some users intend only to hold money offline so that they can transact it online, introduces costs and risks that are entirely avoidable.
+It is possible to build a common framework that supports both online and single-offline settlement involving assets held offline in free hardware devices and dual-offline settlement involving assets held in certified hardware devices.
+Oblivious management offers an opportunity to deliver on this promise by allowing both online service providers and certified hardware devices to serve as integrity providers for digital assets.
 7 Conclusion
 In the end, we must ask ourselves: What problem are we really trying to solve with digital currency?
-Manifestly, the problem is the decline of cash, and the concomitant challenges of maintaining cash infrastructure when the variable revenues of cash infrastructure fall below the fixed costs.
-But the decline of cash really only exists in environments with high network connectivity, wherein dual-offline payments are not necessary.
+Manifestly, the problem is the declining use of cash.
+But this decline is really only present in environments with high network connectivity, wherein dual-offline payments are not necessary.
 For environments with low network connectivity, the decline of cash has not occurred, and cash continues to be the best solution for such environments.
-Many central banks, including the Bank of England, have affirmed a commitment to cash, while underscoring the risks associated with the decline of cash use in favour of online payment schemes such as card payment networks.
-Really, the use cases that digital currency payments should be targeting are cases for which at least one counterparty is already online.
-To argue for digital currencies with the premise that support for dual-offline payments are its principal source of added value mainly supports the view that such products are solutions in search of a problem.
+Many central banks, including the Bank of England, have recently affirmed a commitment to cash, while underscoring the risks associated with the decline of cash use in favour of online payment schemes such as card payment networks.
+This article has argued that the use cases that digital currency payments should be targeting are cases for which at least one counterparty is already online.
+To argue for digital currencies with the premise that support for dual-offline payments are its principal source of added value mainly supports the view that such products are solutions in search of a problem, as the UK House of Lords prominently suggested [House of Lords 2021].
 The purpose of digital wallets (and indeed, digital currency in general) is not primarily about enabling transactions in which both parties are offline.
-The purpose of digital wallets is to allow users to hold assets directly, without a custodian.
-Put differently, this is a way for users to have private property that they can exchange.
+The purpose of digital wallets is to allow users to hold digital monetary bearer instruments directly, without a custodian.
+Put differently, this is a way for users to hold digital assets that they can exchange.
 By facilitating such payments, digital currency can protect the interests of consumers in all cases, including fully networked ones.
-If the counterparty has network connectivity (for example, an electronic point of sale device), then transactions involving digital wallets can be consummated without special hardware.
+If the counterparty has network connectivity (for example, an electronic point of sale device), then transactions involving digital wallets can be consummated without certified hardware.
 Even if we accept that fair exchange is impossible without the participation of a trusted third party, and that the trusted third party can participate via the network or via trusted hardware, it is important to underscore that a non-networked digital wallet can conduct a fair exchange with a networked counterparty, without the aid of special security mechanisms built into the wallet.
@@
 Part of the reason that cash is successful is its role as a bearer instrument.
-Credit systems actually predated cash, and cash emerged as a partial substitute [44].
+Credit systems actually predated cash, and cash emerged as a partial substitute [Graeber 2011].
 What the world does not yet have is an institutionally supported version of cash for the digital economy, and a digital currency aspiring to fill this role will require a non-custodial wallet architecture.
 Custodial wallets require network access as a rule, because consumers must be able to use their devices to access their custodial accounts.
-However, digital wallets do not imply a design choice between network access and certified hardware in the general case.
+However, support for non-custodial wallets does not imply a design choice between network access and certified hardware in the general case.
 In particular, single-offline payments can make use of non-custodial wallets without certified hardware to hold money offline and transact it online, leveraging the network connectivity provided by POS devices presented to customers by merchants.
-To deny the salience of this use of wallets is to impede the development of a technology that may one day become crucial for the digital economy.
+To deny the salience of this use of wallets is to impede the development of a technology that can protect the privacy and autonomy of users while limiting the empowerment of big players in digital payments infrastructure.
 Ultimately, the digital wallets that support economic transactions should be seen not as access devices or proxies offered by custodians, but as tools freely chosen, and even built, by end-users to store and use digital money autonomously, on their own terms.
 Acknowledgements
-The author thanks Professor Tomaso Aste for his continued support for our work on digital currencies and digital payment systems.
+The author thanks Professor Tomaso Aste for his continued support for his ongoing work on digital currencies and digital payment systems.
+The author thanks Christian Grothoff and Yuri Biondi for their generous and insightful feedback on earlier drafts of this article.
 The author also thanks his many colleagues in the international standards development community for a series of vigorous arguments that were instrumental to reaching the crux of this most contentious debate.
-The author also acknowledge the Future of Money Initiative at University College London, the Systemic Risk Centre at the London School of Economics, EPSRC, and the PETRAS Research Centre EP/S035362/1 for the FIRE Project.
+The author also acknowledges the Future of Money Initiative at University College London, the Systemic Risk Centre at the London School of Economics, EPSRC, and the PETRAS Research Centre EP/S035362/1 for the FIRE Project.

### v5

@@
 In principle, retail CBDC would be a direct obligation of the central bank, like cash or central bank reserves, but existing in digital form and held by ordinary consumers and non-financial businesses.
-Some proponents of retail CBDC, including central banks, have argued that it should have cash-like properties to deliver value to end-users [ECB 2021, Bank of England 2023].
+Some proponents of retail CBDC, including central banks, have argued that it should have cash-like properties to deliver value to end-users [ECB 2021, Bank of England and HM Treasury 2023].
 This argument is at least partially right.
@@
 We suggest that the “e-cash” vision of digital currency is largely absent from the discussion of the design of digital currency hardware wallets, and we argue that this neglect is a mistake.
-Contemporary examples of “e-cash” systems with significant deployments include GNU Taler [GNUnet and Taler] and (de facto) Privacy Pass [Davidson 2018], both of which feature the use of non-custodial wallets to hold tokens for redemption.
+Contemporary examples of “e-cash” systems with significant deployments include GNU Taler [GNUnet and Taler] and (de facto) Privacy Pass [Davidson et al.
+2018], both of which feature the use of non-custodial wallets to hold tokens for redemption.
 In both the “offline account” and “e-cash” use cases, dual-offline transactions impose additional security requirements on the transacting parties, in addition to the security requirements applicable to digital transactions in general.
@@
 The asset is stored by the end-user directly and is not recorded on a ledger.
-Double-spending can be prevented in a few ways, one of which is to assume (a) that the issuer processes all transactions as redemptions, (b) that the issuer maintains a ledger to keep track of the set of redeemed assets that it receives in these transactions, and (c) that the issuer, upon receiving an asset in a transaction, checks the ledger to verify that the asset has not already been redeemed [Chaum 1982, Chaum 2021].
+Double-spending can be prevented in a few ways, one of which is to assume (a) that the issuer processes all transactions as redemptions, (b) that the issuer maintains a ledger to keep track of the set of redeemed assets that it receives in these transactions, and (c) that the issuer, upon receiving an asset in a transaction, checks the ledger to verify that the asset has not already been redeemed [Chaum 1983, Chaum et al.
+2021].
 An alternative approach, which we call oblivious management, does not require the issuer to process transactions or maintain a ledger of redeemed tokens.
-In such a system, the assets carry their own provenance information, linking transaction information they carry to rolled-up aggregation entries, one per time period, on a ledger [Goodell 2023, Friolo 2025]; however, neither the tokens nor the value they represent are represented on the ledger at all.
+In such a system, the assets carry their own provenance information, linking transaction information they carry to rolled-up aggregation entries, one per time period, on a ledger [Goodell et al.
+2023, Friolo et al.
+2025]; however, neither the tokens nor the value they represent are represented on the ledger at all.
 In both cases, issuers might keep accounting records of the tokens that they issue or redeem, but the privacy-preserving certification unlinks the issuance of an asset, where the party requesting a new asset may be identified, from its subsequent payment transaction and redemption, wherein only the payee is known.
@@
 ISO 22739:2024 defines a wallet as an “application or mechanism used to generate, manage, store or use private keys and public keys or other digital assets” [ISO 22739 2024a].
-The simplicity of this definition would seem to understate the potentially complex features required to implement token management and custody, an exploration of which is given by Barresi and Zatti [Barresi 2023].
+The simplicity of this definition would seem to understate the potentially complex features required to implement token management and custody, an exploration of which is given by Barresi and Zatti [Barresi and Zatti 2023].
 We address some of the assumptions that Barresi and Zatti make about the design of wallets and suggest different ways of viewing the problem.
-In particular, we critically address the arguments that dual-offline payments constitute the most important use case for digital currency (see Section 2), that trusted computing is an indispensable enabler of digital currency solutions (see Section 3), that digital currency must be held in accounts (see Section 4), and that a wallet is properly seen as “not a place where you can store cash [but] that helps you keep track of it” [Barresi 2023] (see Section 6).
+In particular, we critically address the arguments that dual-offline payments constitute the most important use case for digital currency (see Section 2), that trusted computing is an indispensable enabler of digital currency solutions (see Section 3), that digital currency must be held in accounts (see Section 4), and that a wallet is properly seen as “not a place where you can store cash [but] that helps you keep track of it” [Barresi and Zatti 2023] (see Section 6).
 Additionally, we address the argument implicit to some e-cash solutions that issuance and transaction processing must be performed by the same actor (see Section 5).
-Throughout this article, we shall reference a set of specific designs, including e-cash designs, such as GNU Taler, that are based on bearer instruments, as well as a proposal for how to integrate bearer instruments into a mechanism for using an integrity provider to transfer digital assets obliviously, without maintaining records of the assets, without immediately redeeming them for the value they represent, and without recording any information about the assets or the counterparties involved in transactions [Goodell 2023, Friolo 2025].
+Throughout this article, we shall reference a set of specific designs, including e-cash designs, such as GNU Taler, that are based on bearer instruments, as well as a proposal for how to integrate bearer instruments into a mechanism for using an integrity provider to transfer digital assets obliviously, without maintaining records of the assets, without immediately redeeming them for the value they represent, and without recording any information about the assets or the counterparties involved in transactions [Goodell et al.
+2023, Friolo et al.
+2025].
 The rest of the article is organised as follows.
@@
 2 The sophistry of dual-offline payments
-In addition to the ECB [ECB 2025], several prominent organisations, including the International Monetary Fund [Kiff 2022] as well as other central banks, such as the Bank of England [Bank of England 2022] and the Bank of Canada [2], have released reports suggesting that dual-offline payments are desirable or even essential to the functionality of a future retail CBDC system.
+In addition to the ECB [ECB 2025], several prominent organisations, including the International Monetary Fund [Kiff 2022] as well as other central banks, such as the Bank of England [Bank of England 2022] and the Bank of Canada [Miedema 2020], have released reports suggesting that dual-offline payments are desirable or even essential to the functionality of a future retail CBDC system.
 Such bold statements are striking, considering the prevailing lack of consensus about what the requirements for digital currency should be in general.
@@
 With this in mind, it is reasonable to ask: What problem do proponents of retail CBDC intend to solve, for which dual-offline payments constitute a better solution than online payments involving assets that are held offline?
-We acknowledge that there are some limited real-world use cases for dual-offline payments, such as payments aboard transportation systems that lack reliable network connectivity throughout the system [1].
+We acknowledge that there are some limited real-world use cases for dual-offline payments, such as payments aboard transportation systems that lack reliable network connectivity throughout the system [Clark 2018].
 However, such examples are rare indeed, owing to the prevalence of network connectivity in locations suitable for digital transactions.
-Furthermore, we note that payment scenarios similar to this example do not actually require dual-offline payments; their requirements could also be satisfied by time-shifting, in which a consumer purchases e-money tokens online and consummates transactions with a predetermined merchant offline, by selectively sending to the merchant tokens for which the consumer had already transferred control to the merchant [Goodell 2023].
+Furthermore, we note that payment scenarios similar to this example do not actually require dual-offline payments; their requirements could also be satisfied by time-shifting, in which a consumer purchases e-money tokens online and consummates transactions with a predetermined merchant offline, by selectively sending to the merchant tokens for which the consumer had already transferred control to the merchant [Goodell et al.
+2023].
 (The consumer and the merchant can use a fair exchange protocol to recover any unspent tokens later.) And so, it is fair to ask the question implicit to the title of a report by the UK House of Lords, “Central bank digital currencies: a solution in search of a problem?” [House of Lords 2021].
@@
 However, even if we accept that non-custodial wallets are essential to digital currency, the idea digital currency held offline must be spent offline is certainly false.
-Users have an interest to possess and control digital assets on devices of their choosing, rather than devices associated with attributable accounts, regardless of the context in which they spend them, for reasons of privacy and censorship resistance [Kahn 2005], which we shall address in Section 3.
+Users have an interest to possess and control digital assets on devices of their choosing, rather than devices associated with attributable accounts, regardless of the context in which they spend them, for reasons of privacy and censorship resistance [Kahn et al.
+2005], which we shall address in Section 3.
 Holding digital assets directly takes custodians out of outbound transactions, allowing users to avoid the risk that the custodians might extract fees, block certain (or all) transactions from taking place, or build profiles of the behaviour of their customers by monitoring the details of their transactions, such as time, location, and counterparty.
@@
 Proponents of CBDC are right to acknowledge the secular decline in the use of cash in retail consumer transactions as a primary motivator for digital currency.
-For example, in the UK, banks are closing ATMs and branches [Access to Cash Review 2019], making cash more difficult for consumers to withdraw and for businesses to manage cash [Tisher 2020].
+For example, in the UK, banks are closing ATMs and branches [Access to Cash Review 2019], making cash more difficult for consumers to withdraw and for businesses to manage cash [Tisher et al.
+2020].
 Although court judgments in some jurisdictions suggest that the legal tender status of banknotes implies mandatory acceptance in all but exceptional circumstances [European Union 2021], jurisprudence in other jurisdictions only recognises legal tender as something that must be accepted in the settlement of a debt [Bank of England 2025], admitting the possibility that vendors can freely choose what forms of payment they will accept in the general case.
@@
 The argument that a larger share of domestic transactions should be brought inside the scope of the formal economy is certainly not new [FATF 1996, Camera 2001].
-Recently, the governments of certain countries, including India [Ministry of Finance 2025] and Indonesia [Septiningrum 2024], have taken steps toward eliminating cash entirely, an objective that is consistent with prevailing business-led approaches to phase out cash transactions and encourage digital transactions in general.
+Recently, the governments of certain countries, including India [Ministry of Finance 2025] and Indonesia [Septiningrum et al.
+2024], have taken steps toward eliminating cash entirely, an objective that is consistent with prevailing business-led approaches to phase out cash transactions and encourage digital transactions in general.
 However, achieving that goal, whether it is legitimate or not, does not require CBDC; some countries have already implemented regulations to prevent the use of cash for large transactions or to control bank deposits comprising large amounts of cash [Financial Conduct Authority 2023].
 Moreover, few monetary authorities or government regulators have called for the abolition of cash, whereas many, including those of Switzerland [Reuters 2024], the United Kingdom [Bank of England 2024], and the United States [Federal Reserve 2022], have pledged to preserve access to cash for the foreseeable future.
-Notably, Sweden has rolled back its drive toward creating a cashless society to support cash infrastructure in a near-cashless economy, at great cost [Library of Congress].
+Notably, Sweden has rolled back its drive toward creating a cashless society to support cash infrastructure in a near-cashless economy, at great cost [Library of Congress 2021].
 So the idea that CBDC is mostly a justification for phasing out cash, as has been speculated by some, is not corroborated by the facts, not yet at least.
@@
 In this section, we examine the characteristics of trusted computing within the context of its application to digital currency systems and its potential impact on the power relationships between the users of devices, the manufacturers of devices, and other powerful actors such as corporations and the state.
-Barresi and Zatti explicitly assume the use of trusted hardware to draw transactions into the security envelope of a trusted custodian [Barresi 2023], implicitly drawing the hardware manufacturer into the role of trusted third party.
+Barresi and Zatti explicitly assume the use of trusted hardware to draw transactions into the security envelope of a trusted custodian [Barresi and Zatti 2023], implicitly drawing the hardware manufacturer into the role of trusted third party.
 In their model, the wallet would be a means of transacting with money held in custodial accounts that would be reconciled ex post.
 3.1 Fair exchange
-If we define a fair exchange as a transaction mechanism that ensures that “at the end of an exchange [between two players], either each player receives the item it expects or neither player receives any additional information about the other’s item” [Asokan 1998], then it is not possible to have a fair exchange between two parties without the involvement of a third party; this is a foundational truth [Pagnia 1999].
+If we define a fair exchange as a transaction mechanism that ensures that “at the end of an exchange [between two players], either each player receives the item it expects or neither player receives any additional information about the other’s item” [Asokan 1998], then it is not possible to have a fair exchange between two parties without the involvement of a third party; this is a foundational truth [Pagnia and Gärtner 1999].
 With the exchange of (unregistered) physical objects, possession of the object itself can be sufficient to demonstrate the veracity of the exchange that has taken place.
@@
 To assume that all hardware wallets must take the latter approach because some circumstances might require it is tantamount to overspecification at the expense of users who would readily hold assets offline but transact them online.
-It is worth noting that when Chaum, Grothoff, and Moser argue that “any economically producible device that stores tokens with monetary value in an individual’s possession, and enables offline transactions… will be the target of successful forgery attacks…” [Chaum 2021], they are referring to wallets that perform dual-offline transactions, in contrast to wallets that hold tokens offline and transact them online.
+It is worth noting that when Chaum, Grothoff, and Moser argue that “any economically producible device that stores tokens with monetary value in an individual’s possession, and enables offline transactions… will be the target of successful forgery attacks…” [Chaum et al.
+2021], they are referring to wallets that perform dual-offline transactions, in contrast to wallets that hold tokens offline and transact them online.
 While overspecification of device requirements might certainly appeal to device manufacturers seeking compulsory mandates for their proprietary technology, it is important to consider that protocols that do not rely upon a hardware root of trust can function perfectly well outside the dual-offline context.
@@
 Of course, users with the means to do so can personally conduct the audits themselves or hire contractors to conduct the audits for them.
-However, it is not necessary for every user to actually run an audit; evidence suggests that similar benefits are achieved so long as all users can perform audits if they choose to do so [Heydt-Benjamin 2006].
+However, it is not necessary for every user to actually run an audit; evidence suggests that similar benefits are achieved so long as all users can perform audits if they choose to do so [Heydt-Benjamin et al.
+2006].
   3. 3.
@@
 video streaming).
-Even the specialised Suica card offered by Japan Rail for use at its payment gates, which uses embedded trusted hardware to facilitate payments when connectivity to a central server is broken [Shiibashi 2008], could potentially use locally reachable points of trust to aggregate transaction information rather than assume that network connectivity is either all or nothing.
+Even the use case for the specialised Suica card, which is offered by Japan Rail for use at its payment gates and uses embedded trusted hardware to facilitate payments when connectivity to a central server is broken [Shiibashi 2008], could potentially leverage locally reachable points of trust to aggregate transaction information rather than assume that network connectivity is either all or nothing.
 Moreover, the idea that cash could suddenly become unavailable for general use in offline situations (the latter scenario) is an overly simplistic justification for a system-level requirement.
@@
 A network service that provides payment finality222the state in which obligations are discharged between transacting parties.
-for bearer instruments can be operated by a trusted third party, as in GNU Taler, or it can be oblivious and based on public commitments, so that compromising it does not enable double-spending [Friolo 2025].
+for bearer instruments can be operated by a trusted third party, as in GNU Taler, or it can be oblivious and based on public commitments, so that compromising it does not enable double-spending [Friolo et al.
+2025].
 And so, for single-offline payments, trusted hardware adds nothing, but imposes significant costs and limitations, and therefore it must not be considered a system-level requirement.
@@
 Standards that recommend hardware wallets to use a secure element to prevent the exfiltration of private data might be appropriate in the case of transactions involving smartphones, because of their complexity and exposure to malware attacks.
-However, it is not necessarily the case that all or most digital wallets will be smartphones, despite being the “benchmark in user-accessible features”, as Barresi and Zatti have argued [Barresi 2023].
+However, it is not necessarily the case that all or most digital wallets will be smartphones, despite being the “benchmark in user-accessible features”, as Barresi and Zatti have argued [Barresi and Zatti 2023].
 One reason is price: Although virtually everyone must have a way to engage with the economy, a substantial share of the population cannot afford a smartphone.
@@
 Another reason is security: with ever more complex operating systems and applications, smartphones are increasingly plagued by security vulnerabilities [NIST], many of which specifically target banking credentials [Fadilpašić 2025].
-For a sensitive function like payments, it might be safer and more appropriate to use a dedicated device [Bizama 2025].
+For a sensitive function like payments, it might be safer and more appropriate to use a dedicated device [Bizama et al.
+2025].
 4 Digital currency means tokens, not accounts
-Barresi and Zatti assume that wallets would represent end-user money as account balances [Barresi 2023].
+Barresi and Zatti assume that wallets would represent end-user money as account balances [Barresi and Zatti 2023].
 However, for ordinary consumers and non-financial businesses to hold digital currency directly, there must be some way for end-users to hold digital currency without the requirement to have an attributed account.
@@
 Although the potential for loss or theft of bearer instruments inexorably constitutes a risk that requires a different kind of solution, approaches well-suited to digital technology, such as backup copies and withdrawal records that can be used by victims of theft to selectively de-anonymise their own tokens that had been stolen, can be robust and effective in mitigating such risks.
-Some central banks, including the ECB [European Commission 2023], the US Federal Reserve [Federal Reserve 2022], and the Bank of England [Bank of England 2023], have presented arguments that holding limits for wallets are necessary for financial stability and to prevent bank runs.
+Some central banks, including the ECB [European Commission 2023], the US Federal Reserve [Federal Reserve 2022], and the Bank of England [Bank of England and HM Treasury 2023], have presented arguments that holding limits for wallets are necessary for financial stability and to prevent bank runs.
 However, as an alternative to the heavy-handed approach of monitoring or restricting wallets, an approach of limiting CBDC withdrawals from banks would be no less effective in preventing bank runs or bank disintermediation than limiting cash withdrawals from banks, and the approach of limiting bank withdrawals from banks has been used successfully in the past, notably in Greece in 2015 [Associated Press 2018].
@@
 Importantly, limiting CBDC withdrawals can be done without requiring wallets to implement account-like control mechanisms that undermine the will of their users.
-Barresi and Zatti suggest that wallets could perform “behavioural analytics” on their users [Barresi 2023]; little could be more disrespectful of a user’s autonomy.
+Barresi and Zatti suggest that wallets could perform “behavioural analytics” on their users [Barresi and Zatti 2023]; little could be more disrespectful of a user’s autonomy.
 Similarly, the argument that digital currency could be used as a way to implement “programmable money” that can only be spent in pre-defined ways is a non-starter.
-Such a mechanism would not only undermine the singleness of money but also open a Pandora’s box of ethical concerns [Li 2025], some of which have already been realised by the failed cashless debit card scheme in Australia [Greenacre 2023].
+Such a mechanism would not only undermine the singleness of money but also open a Pandora’s box of ethical concerns [Li and Shen 2025], some of which have already been realised by the failed cashless debit card scheme in Australia [Greenacre 2023].
 However, banks could impose simple rules to limit the aggregate size of withdrawals that can be made from their accounts into CBDC, both in a single transaction and over a period of time, and such rules could be potentially combined with identity information to limit the amount that an end-user could withdraw throughout the system.
-Such rules could also be combined with periodic token expiry, as a mechanism to prevent hoarding [Goodell 2021].
+Such rules could also be combined with periodic token expiry, as a mechanism to prevent hoarding [Goodell et al.
+2021].
 Periodic token expiry is less intrusive than mechanisms that restrict or control wallets directly, and it can be made transparent to users by implementing a mechanism for wallets to automatically refresh any assets that are close to the expiration time.
 Although a third party is necessary for fair exchange (and indeed trusted hardware or closed-source software blobs can act as this third party de facto), Pagnia and Gärtner have demonstrated that the role of the third party can be performed “in zero-knowledge”.
-Specifically, a third party can facilitate final settlement of an asset transfer involving two transacting parties, without having a custodial relationship with either party, without knowing the identity of either party, without knowing how much is being transferred, and even without knowing that tokens are being transferred at all [Pagnia 1999].
+Specifically, a third party can facilitate final settlement of an asset transfer involving two transacting parties, without having a custodial relationship with either party, without knowing the identity of either party, without knowing how much is being transferred, and even without knowing that tokens are being transferred at all [Pagnia and Gärtner 1999].
 Importantly, with oblivious management, an issuer does not rely upon a central ledger to track tokens individually, although the issuer must trust that the system that collects transactions and produces rolled-up aggregation entries, one per time period, will not equivocate.
@@
 Digital assets can indeed be held by users without accounts of any kind, just like cash and other physical assets.
-An example of a digital asset that can be held without an account is an e-cash token in Chaum’s model [Chaum 1982, Chaum 2021], which requires every transaction to function as a redemption, and it has also been shown that digital assets can be transferred without redemption [Goodell 2023, Friolo 2025].
+An example of a digital asset that can be held without an account is an e-cash token in Chaum’s model [Chaum 1983, Chaum et al.
+2021], which requires every transaction to function as a redemption, and it has also been shown that digital assets can be transferred without redemption [Goodell et al.
+2023, Friolo et al.
+2025].
 Accounts for recipients might be desirable, but they are not technically necessary.
@@
 In Chaum’s original proposal, an asset can be transferred to a recipient without revealing information about the payer, and, unlike oblivious management, double-spending is prevented by maintaining a record of the spent tokens.
-In the general case, and in the approach taken by oblivious management, to prevent double-spending, it is sufficient to register non-repudiable commitments from payers, which may take the form of a hash or zero-knowledge proof that can be verified without knowledge of the transaction [Pagnia 1999], and proofs of these commitments can be delivered along with the asset data rather than centrally recorded [Goodell 2023, Friolo 2025].
+In the general case, and in the approach taken by oblivious management, to prevent double-spending, it is sufficient to register non-repudiable commitments from payers, which may take the form of a hash or zero-knowledge proof that can be verified without knowledge of the transaction [Pagnia and Gärtner 1999], and proofs of these commitments can be delivered along with the asset data rather than centrally recorded [Goodell et al.
+2023, Friolo et al.
+2025].
 Oblivious management presupposes that issuers can be held responsible for what they have issued and that proofs of commitments that cryptographically demonstrate the validity and integrity of an asset constitute sufficient evidence of an issuer’s responsibility to honour a redemption.
@@
 Tokens are also important for privacy.
-Systems that make use of transactions involving balances, even if the transactions themselves are privacy-preserving [Wüst 2022, Kiayias 2022], intrinsically require end-users to retain data produced during a transaction that can be used to prove that the transaction was carried out according to the rules.
+Systems that make use of transactions involving balances, even if the transactions themselves are privacy-preserving [Wüst et al.
+2022, Kiayias et al.
+2022], intrinsically require end-users to retain data produced during a transaction that can be used to prove that the transaction was carried out according to the rules.
 This means that transactions are not independent of one another.
-The property of transaction independence is important, because retaining data related to past transactions introduces the risk that the user could be de-anonymised via poor operational security or blackmailed to reveal information about those transactions [Friolo 2025].
+The property of transaction independence is important, because retaining data related to past transactions introduces the risk that the user could be de-anonymised via poor operational security or blackmailed to reveal information about those transactions [Friolo et al.
+2025].
 Because it must be possible for balances to be tracked through all of the movements of that balance, transactions in systems that represent money as balances cannot achieve transaction independence.
@@
 Even if transactions are done with network access to a mutually trusted third party, there is no reason to assume (a) that the third party must be the issuer, or (b) that the issuer must be involved in transaction processing in real-time.
-Chaum’s original proposal assumed both conditions [Chaum 1982], but they are not requirements in general.
+Chaum’s original proposal assumed both conditions [Chaum 1983], but they are not requirements in general.
 In particular, with oblivious management, it is possible to register a (zero-knowledge) commitment to transfer a digital asset without communicating directly with the issuer.
-For a transaction to have finality, assuming every transfer specifies the payee with a unique single-use key, it is sufficient for both parties have confidence in a mutually trusted third party, acting as a notary, to eventually post the commitment (or an aggregation that comprehends the commitment) to an append-only bulletin board that maps single-use keys to commitments and that is trusted by the issuer [Goodell 2023].
+For a transaction to have finality, assuming every transfer specifies the payee with a unique single-use key, it is sufficient for both parties have confidence in a mutually trusted third party, acting as a notary, to eventually post the commitment (or an aggregation that comprehends the commitment) to an append-only bulletin board that maps single-use keys to commitments and that is trusted by the issuer [Goodell et al.
+2023].
 The append-only bulletin board can be operated by the issuer directly, by a party that the issuer trusts, or as a distributed ledger operated by a set of parties that the issuer trusts in aggregate.
-Provided that the operator of the bulletin board does not equivocate, then since no two tokens can be transferred to the same single-use key and have a valid proof of provenance, it is not necessary for the issuer to maintain a ledger that records the registration status of individual tokens [Goodell 2023].
+Provided that the operator of the bulletin board does not equivocate, then since no two tokens can be transferred to the same single-use key and have a valid proof of provenance, it is not necessary for the issuer to maintain a ledger that records the registration status of individual tokens [Goodell et al.
+2023].
 Even if we assume that all tokens are issued by a central issuer, such as a central bank, it is certainly possible to use a decentralised system to conduct transactions using those tokens, and wallets can be designed to respect this possibility.
@@
 If we assume, following Chaum, that individual e-cash assets are generated for one-time use by consumers, then a central bank can provide private-sector banks with “genesis” assets that can be burned, pari passu, by private-sector banks in the process of creating single-use e-cash tokens for consumers.
-Then, the single-use assets can be recirculated within the system by private-sector banks indefinitely via the same process before ultimately being redeemed with the central bank [Friolo 2025].
+Then, the single-use assets can be recirculated within the system by private-sector banks indefinitely via the same process before ultimately being redeemed with the central bank [Friolo et al.
+2025].
 The use of a secure element to assure trust seems to implicitly require the transacting parties to respect the authority of a specific hardware manufacturer, and the power implicit to this authority might be incompatible with the concept of decentralisation if all transacting parties are forced to use secure elements produced by a limited set of manufacturers.
@@
 However, considering the security risks associated with compromised devices, it is better to enforce such rules over the network instead, if it is possible to do so.
-This can be done, for example, by requiring the issuer to validate all transactions, as proposed by Chaum and others [Chaum 1982, Chaum 2021].
-However, there is reason to believe the costs, risks, and public accountability concerns associated with having a single central operator perform issuance, settlement, and compliance functions might be too great for the CBDC use case [Goodell 2021].
+This can be done, for example, by requiring the issuer to validate all transactions, as proposed by Chaum and others [Chaum 1983, Chaum et al.
+2021].
+However, there is reason to believe the costs, risks, and public accountability concerns associated with having a single central operator perform issuance, settlement, and compliance functions might be too great for the CBDC use case [Goodell et al.
+2021].
 A modern alternative to the centralised model involves using a distributed ledger system to put the money itself on a blockchain, the way that UTXO designs (e.g.
@@
 Table 6 provides a comparative overview of the various token-based payment systems that we have mentioned in terms of their key features.
-The digital currency design proposed by Goodell et al [Goodell 2023] and further elaborated by Friolo et al [Friolo 2025] offers an alternative approach to achieving scalability, replacing global consensus about the status and ownership of individual assets, and the related costs of such consensus common to modern cryptocurrency systems, with a combination of local trust and a mechanism for keeping local trust providers honest.
+The digital currency design proposed by Goodell et al [Goodell et al.
+2023] and further elaborated by Friolo et al [Friolo et al.
+2025] offers an alternative approach to achieving scalability, replacing global consensus about the status and ownership of individual assets, and the related costs of such consensus common to modern cryptocurrency systems, with a combination of local trust and a mechanism for keeping local trust providers honest.
 By integrating a fungible token created using Chaum’s method that delivers privacy by design for consumers into a non-fungible asset that can be transferred with strong traceability following the initial creation, it enables the separation of the costs and risks associated with the aforementioned roles that enable payments.
@@
 To address the observed reduction of cash acceptance in urban environments [Mansuri and Vyas 2024, De Nederlandsche Bank 2025] by providing digital currency, whilst encouraging the provision of facilities for converting between cash and digital currency at the boundary of “unbuilt” environments, would be an entirely legitimate approach.
-In particular, if one potential benefit of CBDC in an age of waning cash infrastructure [^4] Evidence for the decline of cash is growing [Access to Cash Review 2019, Tisher 2020], and it is acknowledged as an important risk by central banks, including the ECB [European Commission 2023a].
-is the ability to electronically transport money to remote population centres during crises [Goodell 2020], then there is no reason to require that central bank money distributed in this manner must remain digital.
+In particular, if one potential benefit of CBDC in an age of waning cash infrastructure [^4] Evidence for the decline of cash is growing [Access to Cash Review 2019, Tisher et al.
+2020], and it is acknowledged as an important risk by central banks, including the ECB [European Commission 2023a].
+is the ability to electronically transport money to remote population centres during crises [Goodell et al.
+2020], then there is no reason to require that central bank money distributed in this manner must remain digital.
 For example, service providers in a particular locality can provide physical tokens in exchange for CBDC, where the CBDC is transferred using one of the methods described earlier, much as a casino can provide chips in exchange for cash or bank money.
@@
 Acknowledgements
-The author thanks Professor Tomaso Aste for his continued support for his ongoing work on digital currencies and digital payment systems.
+The author thanks Professor Tomaso Aste for his continued support for our ongoing work on digital currencies and digital payment systems.
 The author thanks Christian Grothoff and Yuri Biondi for their generous and insightful feedback on earlier drafts of this article.